A Bibliography of Publications by, and about, Richard Wesley Hamming

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

10 November 2022 Version 1.16

Abstract

This bibliography records publications of Richard Wesley Hamming.

Title word cross-reference

(2, 2) [KSSY12]. 0/1 [Zie01]. 1 [AIR17, Sal13]. $11 [Jun62]. $12 [Ham68a]. $12.50 [Ham68b]. $17.50 [Ham65a]. 2 [Chu91, vdHL10]. 64 [KY15]. cos X [Ham45b]. d [Siv14]. GF(2) [She95]. H(n, 2) [BBZ17]. k [DG17, SWH12]. m [Tak96]. F_{p}^{N} [Mun06]. Z_{p}^{k} [Mun06]. Mat_{n,s}(Z_{k}) [DS04]. O(n⌈m/w⌉) [GF08]. q [LCL08]. R_{3} [KY15]. t [BBZ17]. τ [WZYG17].

-ary [LCL08]. -Borsuk [Zie01]. -bounded [SWH12]. -D [Sal13]. -designs [BBZ17]. -Dimensional [Siv14]. -lifts [KY15]. -linear [Mun06]. -means [DG17]. -metric [WZYG17]. -nd [Chu91]. -sequences [Tak96].

1983 [Gen83].

2002 [B^{+}02]. 2nd [Blo84, Cun73, S.75].

3 [Bul88]. 3-527-26463-9 [Bul88]. 3-dimensional [BBHS08].

81e [Ham85a]. 87 [Ham85a]. 8th [IS87].

9 [Bul88].

A. [Kar77]. Abstract [BKR01, Tre97]. Academic [Ham64, Ham68d]. access [EQH00]. ACM [ACM11, Ash87]. Addition [Rob87]. Affine [Ser11]. Again [Isl22]. Aids [Ham52, Ham50b]. Algebraic [ORSW12]. Algorithm [Hun93, Rah17]. Algorithms [MS02, SS02, Sti02]. Alphabets [Gil52]. ambiguity [PW14]. AMBTC [KSSY12]. Amer [Ham85a]. America [Dur88]. American [ALW^{+}88]. amplitude [HK78]. Analog [Ham65a, SH99]. analoger [SH99]. Analysis [Ham71d, Ham73c, Ham89c, Ham12a, SH90, SH99, Ham65e, Ham65f, Ham66c, Ham74, Mic85, Sca74, Bra73, Ham65b, Kar77]. Anniversary [Cer80, ALW^{+}88]. annual [ACM97, ACM99]. anticodes [Ahl06]. Aperiodic [ZTYP12, HPL15]. APN [RSV15b]. application [JP13]. Applications [Liu09]. Applied [Bra73, Ham71d, Ham85b, Ham89c, Ham04, Ham12a, Jun62, Kar77, EQH00]. approach [Ham74, Ham91a, RZ12]. Approximability [Tre00, Tre97]. Approximate [KK01, CCGL99, KK03]. arbitrary [SWH12, VR06]. architecture [Chu01]. Arithmetic [IS87, Swa90, Swa15a, Swa15b, Swa15c]. array [Jai07]. Art [Gut92a, Gut92b, Ham63b, Ham91b, Ham95c, Ham97b, MZ02a]. ary [LCL08]. Aspray [Ham92a]. Assignment [MP10]. Associated [MP10]. association [BBZ17]. associative [HT90]. Asymptotic [AJM08]. Atlanta [ACM99]. AtomicHeritage [Kel18a]. Attacks [ORSW12, ZGZ^{+}13]. attributed [Dij76b]. August [B^{+}02]. autocorrelation [PW14]. Automatic [Ham68a]. avoiding [RSV12, RSV15a]. Award [Ash87]. Awarded [Cer80].

baby [Sti02]. baby-step [Sti02]. Back [Isl22]. Bag [WZNM14]. Bag-of-Visual [WZNM14]. Ball [Ham89a]. Balls [LS14]. band [HK78]. band-edge [HK78]. Bandwidth [AJM08, BBHS08, Har03]. Based [Bal96, CTL07, KSSY12, LMWC17, MP10, AB08, AGLP16, CCLL11, HY17, JP13, REF15, ZGZ^{+}13]. Bases [B^{+}02, Sta18, TZZ95, TZZ97]. BCH [Chu91, Sol00]. be [AIR17]. become [ALW^{+}88]. Behavior [PZ06]. behaviour [Ale70]. beholder [Ham94]. Beightler [Ham68b]. Between [Ano14, Kle13]. Beyond [DM97, ORSW12]. Binary [Chu91, HM65, ÖV04, Sta18, Zie01, BBZ17, HT90, KY15]. biological [Swe89]. Bit [EQH00, GF08, Rah17]. Bit-parallel [GF08]. Blum [Ham73c]. Blurring [Ano14]. Book [Bar62, Blo84, Bul88, Gut92a, Gut92b, Ham56a, Ham56b, Ham56c, Ham57a, Ham57b, Ham57c, Ham57d, Ham58a, Ham58b, Ham58c, Ham58d, Ham59a, Ham59b, Ham59c, Ham59d, Ham60, Ham64, Ham65a, Ham65b, Ham68a, Ham68b, Ham68d, Ham73c, Ham90a, Jun62, S.75, Ter78, Ter84, Tho73]. Boolean [Ser11, SWH12]. Bootstrap [GHPS15]. Borsuk [Zie01]. both [AIR17]. Bound [CR12, CCGL99, CR11, Got96, SK10a]. Boundary [AE03, Ham42b, Ham42a]. bounded [AGLP16, SWH12]. Bounds [AGSU15, MP10, HPL15, Jai07, WZYG17, ZLD13]. Brains [Ham51a]. Buchbesprechung [Kno88].

CA [ACM11, Sal13]. Calculation [DM97]. Calculus [Ham66d, Ham68c, Ham84, Ham85b, Ham88c, Ham04]. Campopiano [Jai07]. Campopiano-type [Jai07]. cannot [AIR17]. Capsule [Ham79]. Card [Ehr21]. Care [Rob01]. Case [Kar77, GF08]. Castle [OLGP15]. Cavity [RZ12]. CBI [Cer80]. celebrates [ALW^{+}88]. century [Dur88, MHR80]. certain [OLGP15]. Chain [HS01]. change [HK78]. Channel [ORSW12, ZGZ^{+}13]. Channels [Eli55, Eli56, Eli74, LA01]. chaos [CCLL11]. chaos-and-Hamming [CCLL11]. Character [Ham45b]. Characterization [CVB15, AA05, BvDK08, VR06]. Charles [Ham68b]. Checking [HRSC54, HRSC55]. China [B^{+}02]. Chislennye [Hem68, Khe72]. Chooses [Cer80]. Class [Got96, HP66, Ham71b, SK10b, Ham71a, HY17, TSS04]. Classic [Lew21]. Classical [Ham12b, LL22, MM12]. Classifying [Hun93, Öst02]. Classroom [Ham70b]. Cliffs [Ham68a, Ham68b]. Closing [Gui04]. Clustering [DG17]. cm [Ham64]. cm. [S.75]. Co [S.75]. Code [KSSY12, KS17, Liu09, Sac16, Sta18, ZLD13, CCLL11, Gui04, HPR04, KY15, Rah17]. Codes [Ash98, Bau07, Cer80, Chu91, DS04, EIZ91, EL93, Fen15, GÖ06, Ham50a, Ham90b, Ham95b, Ham12b, Ham21, HH01, HLR01, HS01, LA01, LL22, LWW10, Mit09, MZ02b, PZ06, PL95, Rob87, SK10b, Sch91, Sch03, TZZ95, Zie01, VV94, AB08, CM95, Chu01, GP13, Got96, Gro06, Gui04, Ham09, Ham15a, HY17, HK09, Hou97b, Isa05, KY15, Kle13, LCL08, Mun06, OLGP15, RSV12, RSV15a, RM16, SK10a, TZZ97, TSM14, VR06, WZYG17, Bau07, Ham09, Man16, Sch91, Sch03]. Codierung [Ham87b, Kno88]. Códigos [PP18]. Coding [Eli55, Eli56, Eli74, Gol49, Ham80c, Ham86a, Ham88b, Hou97a, IS86, Khe83, MZ02a, Jai07, MC10, Ham87b]. coincidences [Gro06]. Colleagues [PSvE^{+}89]. Collection [MHR80, Cer80]. College [Ham88c]. Color [Rob01, TH49]. Coloring [Zie01]. combinatoire [CCM92]. combinatorial [CCM92]. Comments [PSvE^{+}89]. Communication [AGSU15, CR12, PE86, CR11, HSZZ06]. Como [IS87, PP18]. Comparing [CDIM02]. Comparison [Gil52]. complement [TSM14]. Completed [Cer80]. completeness [EKP85]. complex [Lev04]. Complexity [AGSU15, CR12, CCGL99, CR11, HL02, HSZZ06, PE86]. Compliant [Ano98a]. Compression [Sal13, SM10, AB08]. Computation [Ham52, Ham73c, Ham50b, JP13, WK14, Ham68a, Ham68a]. Computer [Ano14, Cer80, Ham57e, Ham63a, Ham66d, Ham68c, Ham69a, Ham73b, Ham79, Ham92b, IS87, Lew21, Swa90, Swa15a, Swa15b, Swa15c, Bau09, Ham61b, Ham75b, Ham87a, Mor02, Gen83]. Computers [HRSC54, Ham62a, Ham71c, Ham72, HRSC55, Ham65d, Ham65a, Tho73]. Computing [ACM97, ACM99, ACM11, DM97, Ham63b, HM65, MHR80, SWH12, JP13, MR05, PE86, Ham92a]. concatenated [Isa05]. cone [Avi81]. Conference [B^{+}02, Ham88c, SM10]. Conference/Workshop [Ham88c]. confusions [Ham75b]. conjugates [Sha09]. connection [Kle13]. Connections [Hun93]. Connectivity [FvdHH16]. Constant [Fen15]. constraint [CR02]. constraints [TSM14]. Construction [HPP^{+}17]. constructions [ZLD13]. content [TSM14]. continuous [DG17]. Contrast [HH65]. Contributors [Ham65c]. control [Ham70a]. controlled [Ham69b]. Controlling [Ham57e]. Convergent [Ham45a]. Cornelius [Sca74]. Correcting [Bau07, Chu91, Ham50a, Ham90b, Ham95b, Ham12b, Ham21, Mit09, MZ02a, SK10b, AB08, Got96, HH51, Ham09, Ham15a, LL22]. Correction [Rob87, Chu01, Ham59f, Rah17]. Corrections [Ham48]. corrector [Ham59e]. Correlation [ZTYP12, HPP^{+}17, LPH14, RFZ14]. correlations [HPL15]. Counter [EQH00]. Course [Ham88b]. Created [Lew21]. Creep [HH57, HH58]. Critical [Sta18]. Crossover [MS02]. Crystals [BH53, MH59, MH63]. CS [Ano98a]. Cube [DM95, Sam17, CCGL99, ZGZ^{+}13]. Curriculum [Ham88c]. curve [HK09]. Cut [LCWY12]. cyclic [Chu01].

D [Ham92b, Kar77, Sal13, vdHL10]. Daniel [Kar77]. dans [CCM92]. Data [Ano14, B^{+}02, CDIM02, DG17, Sal13, SM10, AB08]. Databases [Bal96, Bal05]. DDC [SM10]. Dear [Ham66b]. December [Gal09]. Decoder [HH01, KKI^{+}15]. decoding [Sol00]. decomposition [AP12]. defined [KKI^{+}15]. Degenerate [SK10a]. Degree [Cer80]. Delight [War13]. Density [PZ06]. Depend [MS02]. Derivation [TSS04]. Design [LMWC17]. designs [BBZ17]. Detecting [Ham50a, Ham90b, Ham21, HH51, Ham15a]. Detection [Rob87, Ham59f, Rah17]. Determination [AJM08]. Deterministic [AGSU15]. Develop [Ham88c]. diametric [Ahl06]. Did [Ham80b]. Dies [Fis98]. Differences [MS02]. Differential [Ham42b, Ham42a, Kar77, Ham51b, Ham59e]. digit [AW91, AW93]. Digital [BBB17, Fis98, Gol49, HRSC54, Ham57e, Ham77, HS79, HY80, Ham83b, Ham89b, Ham95a, Ham98b, SH90, HRSC55, Cer80, SH99, Blo84, Ter78, Ter84]. Digitale [Ham87c, SH99, Bul88]. Dijkstra [PSvE^{+}89]. Dimension [AE03]. Dimension-Normalized [AE03]. Dimensional [Siv14, AKN13, BBHS08]. Dimensions [Zie01]. discipline [Dij76a]. Discovery [Ham93]. Discrete [Ham84, PW14, Sti02]. Discussion [Ham70b, Ham70c]. Discussions [Ham45b]. Dissimilarity [BKR01]. Distance [AA05, AGSU15, BKR01, CR12, CTL07, DS04, EQH00, GLM02, Ham93, Hun93, IS86, KK01, LCWY12, MP10, PZ06, Sal13, Ser11, Ale70, AIR17, AP12, AGLP16, AD11, CR11, CR02, GF08, HPR04, HSZZ06, Iza09, JP13, KK03, LY08, MR05, PE86, REF15, Sha09, Swe89, TSM14, WK14]. Distances [Bal96, LMWC17, AALS18, Gro06]. distant [Ham98a, Ham01]. Distinction [Ano14]. Distortion [DG17, Ham69b]. Distribution [Ham70d, Ser11, Ham15b]. Distributions [DS04]. Divergent [Ham47, Ham48]. dlya [Hem68, Khe72]. DM [Bul88, Kno88]. DNA [TSS04, TSM14]. do [Ham75a]. document [JP13]. Does [MS02]. Doing [Ham95c, Ham97b]. Dorn [Kar77]. Double [Chu91]. Double-Error [Chu91]. Douglass [Ham68b]. Dr. [Ano14]. Dragonfly [CVB15]. Dual [Chu91, VV94, KY15].

ed [Blo84, Cun73, S.75]. Edge [AJM08, HK78]. Edge-Bandwidth [AJM08]. edit [AALS18]. Edition [Ter84]. Editor [Ham65e]. Edward [Ham64]. Effect [Sta18]. Effectiveness [Isl22, Ham80a, Ham85a]. Effects [Ano14]. Efficient [REF15, SS02, ZGZ^{+}13, KKI^{+}15]. Effort [Ano98a]. Eighth [B^{+}02]. Electron [HH65]. Electronic [Ham65a]. Elementary [Ham70b, Ham70c]. Elements [Ham65b]. Elias [Gal09]. Elliptic [SS02]. Embedding [DG17, WZNM14]. Embeddings [DM95, Mun06, RSV15b]. encoding [Sol00]. encoding/decoding [Sol00]. energy [KKI^{+}15]. energy-efficient [KKI^{+}15]. Engineer [Hou97a]. Engineering [Ham62a, Ham95c, Ham97b, Ham96]. Engineers [Ham62b, Ham73a, Ham86b, Ham91b, Hem68, Khe72, S.75, Bar62, Cun73, Dav63, Gut92a, Gut92b, Jun62]. Englewood [Ham68a, Ham68b]. English [CCM92]. environment [Iza09]. Epstein [Ham68d]. Equations [AKKT16, Ham42b, Ham42a, Ham51b, Ham59e, TSS04, Kar77]. Errata [MH63]. Erratum [RSV15a]. Error [Bau07, Chu91, Ham50a, HH51, Ham90b, Ham95b, Ham09, Ham12b, Ham15a, Ham21, Hou97a, LA01, LL22, Mit09, MZ02a, Rob87, SK10b, AB08, Chu01, Got96, Ham59f, Rah17, RM16]. Error-Correcting [Ham95b, Ham12b, Ham09, LL22, Got96]. Error-detecting [HH51]. Error-Detection [Rob87]. error-reducing [RM16]. erweiterte [Sch91, Sch03]. Essays [HR00, MHR80]. Estendidos [PP18].

Estimating [MP10]. Euclid [Tre97, Tre00]. Evaluation [HH65, HY17, KS17]. Evolutive [GLM02]. exercise [Dij76b]. Existence [ÖV04, ZLD13]. Experience [LL09]. extend [Iza09]. Extended [BKR01, Sch91, Sch03, KY15, Tre97]. Extensions [Rob87]. Extremal [AE03, KY15]. eye [Ham94].

F [Bul88]. Fail [LMWC17]. Fail-Safe [LMWC17]. Fano [Ehr21]. fast [KKI^{+}15]. Fehlerkorrigierende [Bau07, Ham09]. Fejer [Ham92b]. Field [Ham73b]. Fifteenth [Gen83]. fifth [ALW^{+}88]. Fifty [DM97]. filter [HK78, KH77, Ham87c, Bul88]. Filters [Ham77, HS79, HY80, Ham83b, Ham89b, Ham95a, Ham98b, Blo84, Ter78, Ter84]. Finding [Cus13, Ham68b, KK01, KK03]. Finite [Ham68a, HS01]. First [Ham88b, Swe89, ACM99, Ash87]. Fitness [MS02, TSS04]. five [ALW^{+}88]. Fixed [PZ06]. Foreword [Ham90c, Ham96]. Formula [Ham71b]. Formulas [EH72, HP66, Ham71a]. FORTRAN [Kar77]. Forward [Mit09]. Foundations [Ham92b, Ham68b]. Fourier [PW14]. fragile [CCLL11]. Frequency [ZTYP12, Ham74, Ham91a, HPL15, HPP^{+}17, LPH14, RFZ14]. Frequency-Hopping [ZTYP12, RFZ14]. Full [ÖV04]. Full-Rank [ÖV04]. Function [Ser11, HK78]. functionalities [JP13]. Functions [Ham70b, Sam17, Ser11, ZTYP12, Ham70c, SWH12]. Fundamentals [LL22]. Future [Ham83a, Ham90d, HR00, Lew21, Ham00]. Fuzzy [BKR01, Iza09].

G [Ham66a]. Galois [Ash98, Ehr21]. Gambling [Ham68d]. Gap [CR12, CR11]. Gap-Hamming-Distance [CR12, CR11]. Gauss [HY17]. GC [TSM14]. GC-content [TSM14]. Generalized [Ash98, Chu91, GÖ06, HS01, LWW10, VV94, HK09, LCL08, OLGP15, ZLD13]. Genetic [HPR04, MS02]. Geometric [Tre00, Tre97]. Georgia [ACM99]. German [Bau07, Bau09, Ham87b, Ham09, Sch91, Sch03, SH99]. Germany [Bul88]. Giant [Ham51a, Sti02]. giant-step [Sti02]. Glyph [LMWC17]. Golay [Man16, MZ02b, Sol00]. Golay-Codes [Man16]. Granino [Ham65a]. graph [FvdHH16, Har03, vdHL10]. graphes [CCM92]. graphics [KKI^{+}15]. Graphs [AJM08, AE03, Zie01, AA05, AK94, BBHS08, BvDK08, CCM92, GP13, IK97]. Grids [AJM08]. Group [EIZ91, Ahl06].

H [Kar77]. Hacker [War13]. Halfspaces [LS14]. Hall [Ham68a, Ham68b, Ter78, Ter84]. Hammer [Sta18]. Hamming [Bau07, Bul88, IEE18, Jun62, Kno88, RSV15a, S.75, Ter78, Ter84, Ahl06, AA05, AK94, AJM08, AB08, Ale70, AKN13, AGSU15, AIR17, AALS18, Ano98a, Ano98b, Ano04, Ano14, AP12, AGLP16, AW91, AW93, AKKT16, Ash98, AD11, Avi81, AE03, Bal96, Bal05, BBHS08, BvDK08, BBZ17, Bar62, Bau07, Blo84, BBB17, BKR01, Bra73, CVB15, CCM92, Cer80, CCGL99, CR11, CR12, CTL07, CCLL11, CM95, Chu91, Chu01, CDIM02, CR02, Cun73, Cus13, DM95, Dav63, DG17, Dij76b, DS04, Ehr21, EQH00, EIZ91, EL93, EKP85, FvdHH16, Fen15, Fis98, GP13, GM15, Got96, GF08, GHPS15, Gro06, GLM02, GÖ06, Gut92a, Gut92b, Ham90c, Ham93, Ham99, HPL15, HPP^{+}17, Har03, HPR04, HY17, HH01, HK09, HLR01, HL02, HS01, Hou97b, HT90]. Hamming [HSZZ06, Hun93, IK97, Isa05, Isl22, IS86, Iza09, Jai07, JP13, Kai09, KY15, Kar77, KSSY12, KKI^{+}15, Kle13, KS17, KK01, KK03, LA01, LL09, Lee98, LMWC17, Lev04, LY08, LCL08, Liu09, LWW10, LCWY12, LPH14, LS14, LP02, MC10, Man18, MR05, Man16, MS02, Maz03, Mit09, MP10, MZ02b, Mor98, Mor02, Mun06, OR12, OLGP15, ORSW12, Öst02, ÖV04, PP18, PE86, PZ06, Per93, PL95, PW14, Rah17, RZ12, RFZ14, RSV12, RSV15b, REF15, Rob87, RM16, Sac16, SS02, Sal13, Sam17, SK10a, SK10b, Sch91, Sch03, Ser11, Sha09, She95, Siv14, Sol00, Sta18, Sti02, SWH12, Swe89, Tak96, TSS04, Tho73, Tre97, Tre00, TZZ95, TZZ97, TSM14, VR06, WZNM14, WZYG17, WK14, ZGZ^{+}13, ZTYP12, ZLD13, Zie01, VV94, vdHL10]. Hamming- [Man16]. Hamming-autocorrelation [PW14]. Hamming-Codes [Sch91, Sch03]. Hamming-Distance [AGSU15, Ale70]. Handbook [Hou97a]. Harmful [Ano14]. Hashing [Bal96, Bal05]. Heapsort [Kle13]. Henrici [Ham65b]. Hermitian [HK09]. Hiding [CTL07, Ham00]. Hierarchical [KS17]. hierarchy [VR06]. High [Isa05, PZ06, She95]. High-rate [Isa05]. Higher [Rob87]. Hill [Ham65a, Jun62, S.75]. Historical [Bau09]. Historische [Bau09]. History [MHR80]. hit [HPL15, HPP^{+}17, LPH14]. Hong [B^{+}02]. honor [HR00]. Honorary [Cer80]. honour [Sca74]. Hopping [ZTYP12, HPL15, HPP^{+}17, LPH14, RFZ14]. Horse [Sta18]. Houston [Gen83]. Huffman [SK10b]. Human [Ano14]. Humanist [Bra73]. Hybrid [Ham65a]. hyperbolic [Ham51b]. Hypermetric [Avi81].

Ideas [Lew21]. identification [HL02]. Identifying [HLR01]. Illus [Ham65a, Jun62, Ham68a, Ham68b]. Image [CTL07, HH65]. Impact [Ham65d]. Imperfect [TSS04]. Implications [CVB15, Ham63a, Ham61b]. Improvements [GÖ06]. Improving [HT90]. Indexing [WZNM14]. Indirect [Bal96]. Inequality [Sam17]. Infinite [Ham68a]. inflected [HK78]. Informatik [Bau09]. Information [Ham80c, Ham86a, Ham87b, MM12, EQH00, Ham00, Khe83, Ham87b, Kno88]. Informationstechnologie [Bul88]. Informative [WZNM14]. informatsii [Khe83]. Input/Soft [HH01]. Integration [HP66, Ham71b, Ham71a]. Intellectual [Ham61b, Ham63a]. Interface [Gen83]. interleaving [HPP^{+}17]. International [ACM11, B^{+}02]. Intersection [RSV12, RSV15a]. Interview [Kel18b, Kel18a]. intractability [MR05]. Introduction [Ham64, Ham71d, Ham89c, Ham95c, Ham12a, Kar77, Bra73]. Inverse [LY08, LCWY12]. Investigation [Mit09, EQH00]. inzhenerov [Hem68, Khe72]. Irrational [Sta18]. ISBN [Bul88]. Isometric [Mun06]. Italy [IS87]. IV [Kar77]. ix [S.75].

J [Ham68b]. John [Ham66b, Ham92a]. Joint [SS02]. Jose [ACM11]. June [ACM11].

Kendall [WZYG17]. Kernel [WZNM14]. Kernels [PL95]. Know [Ham80b, Ham75a]. kodirovaniya [Khe83]. Kong [B^{+}02]. Konrad [Cer80]. Korn [Ham65a, Ham66a].

L [Ham64, Ham68a]. Lanczos [Sca74]. landscape [TSS04]. Large [B^{+}02]. lattice [AK94]. Launches [Ano98a]. Leakage [ORSW12]. Lean [Ham88c]. Learn [Ham95c, Ham97b]. Learning [Ham95c, Ham97b, Ale70]. Lectures [Ash87]. led [Ham91a]. Length [Fen15, SK10b, KY15]. lesion [TSS04]. Letters [Ham65e]. Level [Ham88c]. lifts [KY15]. Limit [Ser11]. Linear [AKKT16, Ash98, Ham42b, Ham42a, LWW10, HY17, IK97, LCL08, Mun06, VR06]. Lively [Ham88c]. local [REF15]. Location [HM65]. logarithm [Sti02]. Logic [Ham68d]. looking [Cus13]. Loss [Ano14]. lossless [AB08]. Low [DG17, LPH14, Zie01, AKN13, HPL15, HPP^{+}17, Sti02]. low-dimensional [AKN13]. Low-hit-zone [LPH14, HPP^{+}17]. Lower [AGSU15, CR12, CCGL99, CR11, HPL15].

M [Ham65a]. Machine [Ano14, HM65]. Machines [Ham51a, Ham12b, Ham68a]. macromolecules [Swe89]. magnetic [MH59, MH63]. Magnetostriction [BH53]. Man [Ham69a, Man18, Ham87a]. Mandelberg [Kel18a, Kel18b]. Manhattan [MP10]. March [Gen83, SM10]. Martin [Kel18a, Kel18b]. Marvin [Ham68a]. matching [AD11, GF08]. Matchings [AK94]. Math [Ham85a]. Mathematical [Ham47, Ham48, Ham71b, Mic85, Ham92b]. Mathematician [Man18, Mor02]. Mathematicians [Rob01]. Mathematics [Giv62, Ham65e, Ham84, Ham85b, Ham98a, Ham01, Ham04, Isl22, Jun62, Dur88, Ham65f, Ham80a, Ham85a, Ham88a, Ham64]. Matrices [MP10, HPR04]. Matrix [VR06]. max [LY08]. Maximum [DS04]. May [ACM97, ACM99, IS87]. McCracken [Kar77]. McGraw [Ham65a, Jun62, S.75]. McGraw-Hill [Ham65a, Jun62, S.75]. mean [Swe89]. means [DG17]. Measure [BKR01]. Measurement [BH53, Ham73b]. Measurements [Ham66a]. measures [AGLP16]. Measuring [TH49]. Mechanization [Ham61a]. Medal [IEE18]. Meets [Bra73, SK10b, Tre00, Tre97]. Melas [VV94]. Memory [LA01, CM95, HT90]. Mentor [Man18]. Methods [Bar62, Cun73, Dav63, Ham62b, Ham73a, Ham85b, Ham86b, Ham88c, Ham04, Jun62, Kar77, S.75, EQH00, Ham51b, Ham59e, Hem68, Khe72, Kle13]. metody [Hem68, Khe72]. Metric [Bal05, DG17, DS04, AKN13, WZYG17]. Metropolis [HR00]. Microcircuit [Ale70]. min [LY08]. minimal [AW91, AW93]. Minimizing [AE03]. Minimum [LCWY12, PZ06, AKN13]. Minsky [Ham68a]. Mission [Sta18].

Mission-Critical [Sta18]. Model [ORSW12, Swe89, TSS04]. modeling [REF15]. Modern [Ham70a, Ham92a, LL22]. moments [SWH12, Swe89]. Monotone [Ham45a, Ham46, Ham47, Ham48]. monotonicity [AA05]. Monthly [Ham85a]. MR0559142 [Ham85a]. MST [Tre97]. Muller [MZ02b]. Multidimensional [AJM08]. Multiple [MC10, KH77]. Multiplication [SS02]. Multisection [HK78]. Museum [Cer80]. mutation [Swe89]. my [Ham75b].

N.J [Ham68a, Ham68b]. Nature [Ham70b, Ham70c]. nauchnykh [Hem68, Khe72]. nd [Chu91]. nearest [CCGL99]. nearest-neighbor [CCGL99]. neighbor [CCGL99]. Neighbour [GP13]. nets [Ale70]. Network [Chu01, Hun93, Maz03]. Networks [CVB15, DM95]. Neumann [Ham92a]. Neural [Maz03]. News [Cer80]. Next [DM97]. Nick [HR00]. ninth [ACM97]. no [Ham85a]. Node [Hun93]. noise [TH49]. Noisy [Eli55, Eli56, Eli74]. Nominal [DG17]. Nominal-continuous [DG17]. Non [DS04, Jai07]. Non-Hamming [DS04, Jai07]. Noninterpolatory [EH72]. Nonlinear [PL95]. nonrecursive [KH77]. Norm [ZTYP12]. Normalized [AE03]. Norms [CDIM02]. Note [Fen15, HM65, HH58]. Notes [Gol49, Ham45b, Ham47, Ham48, Ham70b, Ham71b, Bau09]. Notices [Cer80]. Notizen [Bau09]. Novel [CTL07, AB08]. November [Gal09]. NP [EKP85]. NP-completeness [EKP85]. Nuclear [MH59, MH63]. Number [Hun93]. Numbers [Ham70d, Ham15b]. Numeral [Sta18]. Numerical [Bra73, Ham62b, Ham64, Ham65b, Ham65e, Ham65f, Ham71d, Ham73a, Ham86b, Ham89c, Ham12a, HH65, Kar77, Ham66c, Ham74, Sca74, Hem68, Khe72, S.75, Bar62, Cun73, Dav63, Ham73c, Jun62].

Observations [Bal96]. Olmo [IS87]. One [Ham69a, Ham87a]. Opening [Cer80]. Optima [Ham68b]. Optimal [CR12, DM95, SK10b, Zie01, Ahl06, CR11, HPP^{+}17, LPH14, RFZ14]. Optimization [Ham68b]. order [RSV15b, She95]. Ordinary [Kar77, Ham59e]. Origins [Ham92a]. Other [Ham50b, Ham52, Kle13]. Output [HH01].

P [Ham92b, Ham64]. packing [RZ12]. Pairs [Gro06]. Paper [HH58]. Papers [Lew21, Sca74]. parabolic [Ham51b]. Paradox [Ham89a, Sta18]. Parallel [KS17, GF08]. Parameterized [AKKT16]. Part [Ham95a]. partial [Ham51b, HPP^{+}17, LPH14]. Paso [ACM97]. Passage [Ano98b]. Pattern [AD11]. peak [TSS04]. People [Ham12b]. Percolation [Siv14, GHPS15]. perfect [Gui04]. Perfeitos [PP18]. Performance [EQH00]. Period [AALS18]. Permanent [Cer80]. permutation [WZYG17]. permutations [RSV15b]. Perspectives [Ham79]. Peter [Ham65b, Gal09]. Phase [HH65, vdHL10]. philosophy [Ham75b]. Phylogenetic [AP12]. physical [Mic85]. physics [Ham00]. Pioneer [Fis98]. planet [Ham98a, Ham01]. Point [HM65, HK09]. Polya [HPS98]. polynomial [SWH12]. Polynomials [She95, Lev04]. Pooling [Ano14]. post [TSM14]. post-processing [TSM14]. pp [Bul88, Ham65a, Ham68a, Ham68b, Jun62, S.75, Ter78, Ter84]. pp. [Ham68d]. Practice [Ham73c]. pre [TSM14]. pre-processing [TSM14]. Predicted [Ano14]. predictor [Ham59e]. Preis [Kno88]. prejudices [Ham75b]. Prentice [Ham68a, Ham68b, Ter78, Ter84]. Prentice-Hall [Ham68a, Ham68b, Ter78, Ter84]. prescribed [Gro06]. PRESENT [ZGZ^{+}13]. Press [Ham64, Ham68d]. Price [Bul88]. Primitive [She95]. Privacy [Ano14]. private [WK14]. Probability [Gut92a, Gut92b, Ham85b, Ham91b, Ham04, LA01, MS02]. Problem [LCWY12, Zie01, AKN13, CCM92, EKP85, HL02, HSZZ06, LY08, Sti02]. problème [CCM92]. Problems [AGSU15, EDD^{+}96, Ham42b, Ham42a, HPS98, MP10, CR02]. Proceedings [ACM97, ACM99, ACM11, IS87, B^{+}02, Gen83, SM10]. Process [Ham66a]. processing [KKI^{+}15, TSM14]. Programming [MP10, Dij76a]. Properties [Bal05, LPH14]. pseudo [SWH12]. pseudo-Boolean [SWH12]. pseudorandom [PW14]. Publications [Dav63, Ham51a]. Pylyshyn [Ham79].

Quadrance [DG17]. Quadratic [MP10]. Quadrature [EH72, Kar77]. Quantum [AGSU15, MM12, Got96, SK10a]. Quasi [LMWC17]. Quasi-Hamming [LMWC17]. quasiperiodic [AIR17]. quasispecies [TSS04].

R [Bar62, Blo84, Bra73, Bul88, Cun73, Dav63, Kno88, Ter84]. R. [Dij76b, S.75]. rabotnikov [Hem68, Khe72]. radio [KKI^{+}15]. radius [SWH12]. Radix [Rob87]. Random [IS86, Ser11, vdHL10, FvdHH16, Ham66a]. Random-Process [Ham66a]. Rank [ÖV04]. rate [Isa05]. reasons [ALW^{+}88]. Recognizing [IK97]. recovery [AALS18]. Reduced [SS02]. Reduces [Hun93]. reducing [Ham69b, RM16]. Redundancy [HRSC54, HRSC55, Chu01]. Reed [Mit09, MZ02b]. Relative [BBZ17, LWW10, LCL08, ZLD13]. Relaxation [HH57, HH58]. relevant [Ham75a]. Remote [Ham69b]. repair [TSS04]. Repeats [GLM02]. Repetitions [KK01, KK03]. replication [Swe89]. replication-mutation [Swe89]. Report [Ham88c]. Reports [Giv62]. Representation [SS02, REF15]. representations [AW91, AW93]. Research [Ham86c, Kai09]. Resolving [ÖV04]. resonance [MH59, MH63]. Respond [PSvE^{+}89]. response [KH77]. responses [HK78]. retrieval [EQH00]. reverse [TSM14]. reverse-complement [TSM14]. Review [Bar62, Blo84, Bul88, Cun73, Gut92a, Gut92b, Ham64, Ham65a, Ham65b, Ham66a, Ham68a, Ham68b, Ham68d, Ham73c, Ham90a, Ham92a, Ham92b, Jun62, S.75, Ter78, Ter84, Tho73, Bra73]. Reviews [Ham56a, Ham56b, Ham56c, Ham57a, Ham57b, Ham57c, Ham57d, Ham58a, Ham58b, Ham58c, Ham58d, Ham59a, Ham59b, Ham59c, Ham59d, Ham60, Ham79, Ham88b, Ham88c, Kar77]. Revolution [Ham63a, Ham66d, Ham68c, Ham61b, Ham79]. Richard [Ano14, Gut92a, Gut92b, IEE18, Jun62, Mor02, Ter78, Tho73, Ano98a, Ano98b, Ano04, Bau07, Fis98, Kai09, Kar77, Lee98, LP02, Man18, Mor98, OR12, Per93, Sac16]. Ring [Ash98]. Rings [HS01]. Role [Ham73b]. Routing [CVB15]. Row [Sta18]. Russian [Hem68, Khe72, Khe83].

S [Ham68b, Kar77]. sa [CCM92]. Safe [LMWC17]. salesman [EKP85]. saliency [REF15]. same [KH77]. San [ACM11]. SAR [B^{+}02]. satisfaction [CR02]. satisfying [TSM14]. saturating [Got96]. Scalar [SS02]. Scalars [SS02]. Scheme [CTL07, HPS98, KSSY12, LMWC17, AB08, BBZ17, CCLL11, GM15]. Science [Ham61a, Ham69a, Ham94, Ham95c, Ham97b, Lew21, Bau09, Gen83, Ham75b, Ham87a, Ham92b]. Scientific [Ham63b]. scientist [ALW^{+}88, Mor02, ALW^{+}88]. Scientists [Bar62, Cun73, Dav63, Gut92a, Gut92b, Ham62b, Ham73a, Ham86b, Ham91b, Jun62, S.75, Hem68, Khe72]. Scratchpad [CCM92]. searching [CCGL99]. Second [Ter84, HK09, OLGP15, Swe89]. Secret [KSSY12]. Section [Giv62]. Sections [Giv62]. Secure [JP13, CCLL11]. Seiten [Kno88]. Self [RSV15b, KY15]. self-dual [KY15]. Self-embeddings [RSV15b]. Semantic [WZNM14]. semiconservative [TSS04]. Semidefinite [MP10]. Sequence [AGLP16, LPH14]. Sequences [ZTYP12, HPL15, HPP^{+}17, PW14, RFZ14, Tak96]. Sequential [IS86]. serially [Isa05]. Series [Ham45a, Ham47, Ham48, Ham68a, Ham46]. Set [Ser11]. Sets [AE03, HLR01, LPH14, RFZ14]. Seventy [ALW^{+}88]. seventy-fifth [ALW^{+}88]. Seventy-five [ALW^{+}88]. Sharing [KSSY12]. sharpen [HK78]. Sharpening [KH77]. Shortened [LA01]. should [Ham12b]. Side [ORSW12, ZGZ^{+}13]. Side-Channel [ORSW12, ZGZ^{+}13]. Signal [SH90, SH99]. Signale [SH99]. Signalling [Gil52]. Signed [AW91, AW93]. similarity [AGLP16, JP13]. Simovici [Ham92b]. Simple [HH01]. Simulation [Ham62a, Ham66a, Maz03, Ham75a]. Simultaneous [SS02]. Single [BH53, TSS04]. single-fitness [TSS04]. Site [Cer80, Siv14]. Slow [Ano14]. Small [HRSC54, TZZ95, TZZ97, HRSC55, RSV15b]. Snowbird [SM10]. Societies [Giv62, Ham73b]. Society [Ham71c, Ham72, Tho73]. Soft [HH01]. Soft-Input [HH01]. Soft-Input/Soft-Output [HH01]. Soft-Output [HH01]. Software [Ham96, KKI^{+}15]. software-defined [KKI^{+}15]. sojourn [Tak96]. Solomon [Mit09]. Solution [Kar77, CCM92, TSS04]. Solutions [EDD^{+}96, HPS98]. Solving [AKKT16, Ham51b]. Some [GÖ06, Ham42b, Ham42a, Ham51b, Sti02]. Sources [MC10]. Space [Bal05, DG17, EL93, IK97, Mun06, RZ12]. Spaces [EIZ91, HLR01, Öst02, ÖV04, Ahl06, Avi81, HL02, Lev04]. spanning [LY08]. sparse [REF15]. Spectral [BvDK08]. Sphere [DG17, RZ12]. spheres [SWH12]. Splitting [Hun93]. Stable [Ham59e]. Standards [Ano98a]. State [Ham63b]. Statistical [Ham68d]. Statistics [Bra73, Ham66c, Ham85b, Ham90d, Ham04, Gen83]. Steiner [AKN13, RSV15b, Tre00]. step [Sti02]. Stiefel [Ham64]. STOC’11 [ACM11]. stochastic [HPR04, Swe89]. Strachey [Cer80]. Streams [CDIM02]. string [GF08]. strings [AIR17, AALS18]. Stroud [Kar77]. Studies [Sca74, Kar77]. Subcodes [LWW10, LCL08, RSV12, RSV15a]. subgraphs [FvdHH16, vdHL10]. Subseries [Ham47, Ham48]. Subspaces [Öst02]. Subtraction [Rob87]. Sum [LCWY12, LY08]. Sum-Type [LCWY12, LY08]. sums [HY17]. supercritical [vdHL10]. Superimposed [EL93]. symmetric [KH77]. Symposium [ACM97, ACM99, ACM11, Gen83, Ham59f, IS87]. System [Sta18, HH51, Ham69b, Lev04]. Systems [Ham62a, Sta18, Mic85, RSV15b]. Systolic [Maz03].

tables [Cus13]. Tandem [GLM02]. Teaching [Ham88c]. Technical [Ham73b]. techniques [HPP^{+}17]. Technology [Ano14, Fis98]. Tennis [Ham89a]. Teoriya [Khe83]. test [Tak96]. tests [AP12]. Texas [ACM97, Gen83]. Their [Chu91, Rob01, LCL08, LWW10]. theorem [Ahl06]. Theory [ACM97, ACM99, ACM11, Ham42b, Ham42a, Ham68d, Ham73c, Ham80c, Ham86a, Ham87b, Ham70a, Khe83, Ham88b]. Theresa [Ham65a]. Thermodynamic [TSM14]. Think [Ham51a, Ham97a, Ham12b]. Thinking [Ano14]. thirty [ACM99]. thirty-first [ACM99]. Thought [Ham80b]. threshold [FvdHH16]. thresholds [AD11]. Tilings [ÖV04]. time [GF08, IK97, SWH12, Tak96]. Topological [CVB15]. TOPSIS [Iza09].

Torus [Siv14, GHPS15]. Trace [GÖ06].

Transcendental

[Ham45b, Ham70b, Ham70c]. transform [PW14]. transitivity [GP13]. transmission [Chu01]. Tree [Tre00, AKN13, LY08]. trees [AP12]. Trends [Ham97a]. Tribute [Ano98a]. Trick [Ehr21]. triple [RSV15b]. Trojan [Sta18]. truth [Cus13]. TSP [Tre97, Tre00]. turbo [CM95, Gui04, Mit09]. Turing [Ash87]. Twentieth [MHR80]. Twenty [B^{+}02, ACM97, Ash87]. Twenty-Eighth [B^{+}02]. twenty-ninth [ACM97]. Two [AIR17, Eli56, Liu09, HK09]. two-point [HK09]. Type [LCWY12, Jai07, LY08].

Undetected [LA01]. Uniform [DS04]. Unit [CM95]. Unit-memory [CM95]. units [KKI^{+}15]. Unreasonable [Isl22, Ham80a, Ham85a]. untitled [Ham90a]. Update [Ano98a]. Urn [HPS98]. USA [ACM11]. Use [Bal05, Rob87, Ham88a, KH77, REF15]. Using [CDIM02, GLM02, Iza09, HK78, Isa05, KKI^{+}15]. Utah [SM10].

Value [Ham42b, Ham42a]. Variable [Fen15, SK10b]. Variable-Length [Fen15, SK10b]. VCH [Bul88, Kno88]. Verarbeitung [SH99]. Verlagsgesellschaft [Bul88, Kno88]. versus [TSM14]. Vertical [EQH00]. Vertices [HLR01]. Very [B^{+}02]. via [AP12, Sol00]. Video [WZNM14]. View [Ham69a, Ham87a]. Villa [IS87]. Visual [WZNM14, REF15]. Visualization [LMWC17]. VLDB [B^{+}02]. Volume [Ham92b]. Vorwort [Ham99]. VQ [CTL07]. vs [Ham65e, Ham65f].

W [Ano98b, Ano14, Bar62, Blo84, Bra73, Bul88, Cun73, Dav63, Dij76b, Gut92a, Gut92b, Ham79, Ham92a, IEE18, Jun62, Kar77, Kno88, LP02, Per93, S.75, Ter78, Ter84, Tho73]. was [Ham91a]. watermarking [CCLL11]. Weight [AKKT16, Chu91, LS14, ORSW12, SS02, She95, TZZ95, AW91, AW93, HK09, LCL08, OLGP15, Sti02, Tak96, TZZ97, VR06, ZGZ^{+}13, ZLD13]. weight-based [ZGZ^{+}13]. Weighted [LCWY12, LY08]. Weights [Ash98, Fen15, GÖ06, HS01, LWW10, VV94, Cus13, HY17]. Weinheim [Bul88, Kno88]. Wesley [Lee98, Man18, Mor98, OR12, Mor02]. Wigner [Isl22]. Wilde [Ham68b]. William [Kar77]. Window [BBB17]. without [Cus13]. Words [WZNM14]. work [Ham12b]. Workshop [Ham88c]. World [BBB17]. worst [GF08]. Would [Ham80b].

x [Ham64]. xiv [Ham68b]. xvi [Ham68d]. xviii [Ham68a]. xxiv [Ham65a].

Y2K [Ano98a]. Y2K-Compliant [Ano98a]. Years [DM97, Ash87]. York [Ham64, Ham65a, Ham68d, Jun62, S.75].

Z. [Ham79]. Zero [CDIM02]. zone [HPL15, HPP^{+}17, LPH14]. zur [Bau09]. Zuse [Cer80].

References

Aider:2005:DMN

[AA05] Méziane Aïder and Mustapha Aouchiche. Distance monotonicity and a new characterization of Hamming graphs. Information Processing Letters, 96(6):207–213, December 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amir:2018:PRS

[AALS18] Amihood Amir, Mika Amit, Gad M. Landau, and Dina Sokol. Period recovery of strings over the Hamming and edit distances. Theoretical Computer Science, 710(??): 2–18, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/S0304397517307752.

Al-Bahadili:2008:NLD

[AB08] Hussein Al-Bahadili. A novel lossless data compression scheme based on the error correcting Hamming codes. Computers and Mathematics with Applications, 56(1): 143–150, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/S0898122107008334.

ACM:1997:PTN

[ACM97] ACM, editor. Proceedings of the twenty-ninth annual ACM Symposium on the Theory of Computing: El Paso, Texas, May 4–6, 1997. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997. ACM order no. 508970.

ACM:1999:PTF

[ACM99] ACM, editor. Proceedings of the thirty-first annual ACM Symposium on Theory of Computing: Atlanta, Georgia, May 1–4, 1999. ACM Press, New York, NY 10036, USA, 1999. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999. ACM order number 508990.

ACM:2011:SPA

[ACM11] ACM, editor. STOC’11: Proceedings of the 2011 ACM International Symposium on Theory of Computing: June 6–8, 2011, San Jose, CA, USA. ACM Press, New York, NY 10036, USA, 2011. ISBN 1-4503-0691-8. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Atallah:2011:PMH

[AD11] Mikhail J. Atallah and Timothy W. Duket. Pattern matching in the Hamming distance with thresholds. Information Processing Letters, 111(14):674–677, July 31, 2011. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/S0020019011001049.

Azizouglu:2003:ESM

[AE03] M. Cemil Azizouglu and Ömer Eugeciouglu. Extremal sets minimizing dimension-normalized boundary in Hamming graphs. SIAM Journal on Discrete Mathematics, 17(2):219– 236, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/37505.

Apostolico:2016:SSM

[AGLP16] Alberto Apostolico, Concettina Guerra, Gad M. Landau, and Cinzia Pizzi. Sequence similarity measures based on bounded Hamming distance. Theoretical Computer Science, 638(??):76–90, July 25, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/S0304397516000475.

Ambainis:2015:LBD

[AGSU15] Andris Ambainis, William Gasarch, Aravind Srinivasan, and Andrey Utis. Lower bounds on the deterministic and quantum communication complexity of Hammingdistance problems. ACM Transactions on Computation Theory, 7(3): 10:1–10:??, July 2015. CODEN ???? ISSN 1942-3454 (print), 1942-3462 (electronic).

Ahlswede:2006:ADT

[Ahl06] R. Ahlswede. Another diametric theorem in Hamming spaces: optimal group anticodes. In 2006 IEEE Information Theory Workshop. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006.

Amir:2017:TSH

[AIR17] Amihood Amir, Costas S. Iliopoulos, and Jakub Radoszewski. Two strings at Hamming distance 1 cannot be both quasiperiodic. Information Processing Letters, 128(??): 54–57, December 2017. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.com/
                                        science/article/pii/S0020019017301473.

Akhtar:2008:ADE

[AJM08] Reza Akhtar, Tao Jiang, and Zevi Miller. Asymptotic determination of edge-bandwidth of multidimensional grids and Hamming graphs. SIAM Journal on Discrete Mathematics, 22(2):425–449, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aigner:1994:MLG

[AK94] Martin Aigner and Regina Klimmek. Matchings in lattice graphs and Hamming graphs. Combinatorics, Probability and Computing, 3(2):157–166, June 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Arvind:2016:SLE

[AKKT16] V. Arvind, Johannes Köbler, Sebastian Kuhnert, and Jacobo Torán. Solving linear equations parameterized by Hamming weight. Algorithmica, 75(2):322–338, June 2016. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic).  URL http://link.
springer.com/article/10.1007/
s00453-015-0098-3.

Althaus:2013:LDS

[AKN13] Ernst Althaus, Joschka Kupilas, and Rouven Naujoks. On the lowdimensional Steiner minimum tree problem in Hamming metric. Theoretical Computer Science, 505(??): 2–10, September 23, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/S0304397513001205.

Aleksander:1970:MLN

[Ale70] I. Aleksander. Microcircuit learning nets: Hamming-distance behaviour. Electronics Letters, 6(5): 134–??, 1970. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Arnheim:1988:SFR

[ALW^{+}88] Rudolf Arnheim, Thomas E. Lovejoy, David Gordon Wilson, Freeman Dyson, Jane Goodall, Ian Shelton, Kenneth H. Olsen, Irene C. Peden, Richard W. Hamming, Thomas Eisner, Preston Cloud, Matt Cartmill, Samuel C. Florman, Jeremy Bernstein, George A. Miller, Robert M. May, G. Evelyn Hutchinson, Jerome Bruner, Priscilla C. Grew, William Bevan, Elisabeth S. Vrba, Myrdene Anderson, Kevin Padian, Harry Shipman, Victor F. Weisskopf, Walter A. Hill, Patricia D. Moehlman, Melvin Kranzberg, Malak Kotb, Raymond Kurzweil, Marcia Mc-Nutt, Masakazu Konishi, Miriam Rothschild, Edward Teller, Alison Jolly, H. Jane Brockmann, Keith Stewart Thomson, Peter J. Denning, Benoît B. Mandelbrot, Abraham Pais, Paul MacCready, Kip S. Thorne, Ruth Sager, Gerald J. Wasserburg, Neal E. Miller, Rita Levi-Montalcini, Stephen Jay Gould, Edwin H. Land, Michel Boudart, Anne Kernan, Douglas R. Hofstadter, Rosalyn S. Yalow, Bruce H. Tiffney, Mimi Koehl, Walter E. Massey, David P. Billington, John A. W. Kirsch, Abner Shimony, J. Donald Fernie, Brian J. Skinner, Lynn Margulis, Sheldon Lee Glashow, Michael LaBarbera, J. Tuzo Wilson, E. R. Ward Neale, Rudolf Peierls, Roald Hoffmann, Mary L. Good, Donald R. Griffin, Vaclav Smil, Michael S. Turner, Sarah Ann Woodin, Luis Alvarez, George A. Bartholomew, and George B. Schaller. Seventyfive reasons to become a scientist:

American Scientist celebrates its seventy-fifth anniversary. American Scientist, 76(5):450–463, September 1988. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.

jstor.org/stable/27855384;
http://www.jstor.org/stable/
pdfplus/27855384.pdf.

Anonymous:1998:CUCa

[Ano98a] Anonymous. CS update: CS launches Y2K-compliant standards effort; tribute to Richard Hamming. Computer, 31(3):85–87, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://pdf.

computer.org/co/books/co1998/
pdf/r3085.pdf.

Anonymous:1998:PRW

[Ano98b] Anonymous. Passage: Richard W. Hamming, 82. Wired, ?? (??):??, January 8, 1998. CO-DEN WREDEM. ISSN 1059-1028 (print), 1078-3148 (elec-

tronic). URL https://www.wired.
com/1998/01/passage-richard-
w-hamming-82/.

Anonymous:2004:RH

[Ano04] Anonymous. Richard Hamming.

In Simon Read, editor, Biographies of Famous Computer Scientists, page 24. Creative Commons, ????,

2004. URL https://archive.org/
details/ost-computer-science-
biographies_screen. From the

introduction: “The biographies in this document have been researched and written by students as an assignment in an Introduction to Computer Science course and then checked and edited by the editor of this document”.

Anonymous:2014:DRW

[Ano14] Anonymous. In 1961 Dr. Richard W. Hamming predicted the harmful effects of computer technology: Loss of privacy, pooling of data and the slow blurring of the distinction between human and machine “thinking”. Web site., December 13, 2014. URL http:

                                        //stuffnobodycaresabout.com/
                                        2014/12/13/harmful-effects-computer-
                                        technology-predicted-1961/.

Anselmo:2012:PTH

[AP12] Cézar A. F. Anselmo and Aluísio Pinheiro. Phylogenetic trees via Hamming distance decomposition tests. Journal of Statistical Computation and Simulation, 82(9):1287– 1297, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ashenhurst:1987:ATA

[Ash87] Robert L. Ashenhurst, editor. ACM Turing Award Lectures: the first twenty years, 1966–1985. ACM Press anthology series. ACM Press and Addison-Wesley, New York, NY 10036, USA and Reading, MA, USA, 1987. ISBN 0-201-07794-9. xviii + 483 pp. LCCN QA76.24 .A33 1987.

Ashikhmin:1998:GHW

[Ash98] Alexei Ashikhmin. On generalized Hamming weights for Galois ring linear codes. Designs, Codes, and Cryptography, 14(2):107–126, May 1998. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (elec-

tronic). URL http://www.wkap.
nl/oasis.htm/156974.

Avis:1981:HSH

[Avi81] David Avis. Hypermetric spaces and the Hamming cone. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33(??):795– 802, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arno:1991:SDR

[AW91] Steven Arno and Ferrell S. Wheeler.

Signed digit representations of minimal Hamming weight. Technical report SRC-TR-91-046, Supercomputing Research Center: IDA, Lanham, MD, USA, July 1991. 18 pp.

Arno:1993:SDR

[AW93] S. Arno and F. S. Wheeler. Signed digit representations of minimal Hamming weight. IEEE Transactions on Computers, 42(8):1007– 1010, August 1993. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=238495.

Bernstein:2002:VPT

[B^{+}02] Philip A. Bernstein et al., editors.

VLDB 2002: proceedings of the Twenty-Eighth International Conference on Very Large Data Bases, Hong Kong SAR, China, 20–23 August 2002. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2002. ISBN 1-55860-869-9. LCCN ????

Balakirsky:1996:HDB

[Bal96] V. B. Balakirsky. Hashing of databases based on indirect observations of Hamming distances. IEEE Transactions on Information Theory, 42(2):664–671, March 1, 1996. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Balakirsky:2005:HDU

[Bal05] Vladimir B. Balakirsky. Hashing of databases with the use of metric properties of the Hamming space. The Computer Journal, 48(1):4–16, January 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.uk/
                                        computer_journal/hdb/Volume_
                                        48/Issue_01/bxh059.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_48/Issue_
                                        01/pdf/bxh059.pdf.

Barnett:1962:BRB

[Bar62] V. D. Barnett. Book review: Numerical Methods for Scientists and Engineers, by R. W. Hamming. Journal of the Royal Statistical Society. Series A (General), 125(4):642– 643, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982636.

Bauer:2007:RHF

[Bau07] Friedrich L. Bauer. Richard Hamming: Fehlerkorrigierende Codes. (German) [Richard Hamming: Error correcting codes]. Informatik Spektrum, 30(2):??, ???? 2007. CO-DEN INSKDW. URL http://

                                        link.springer.com/article/10.
                                        1007/s00287-006-0135-3.

Bauer:2009:RHF

[Bau09] Friedrich L. Bauer. Historische Notizen zur Informatik. (German) [Historical notes on computer science]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-85790-7. x + 454 pp. LCCN QA76.17 .B38 2009.

Bojkovic:2017:HWD

[BBB17] Zoran S. Bojkovic, Bojan M. Bakmaz, and Miodrag R. Bakmaz. Hamming window to the digital world. Proceedings of the IEEE, 105(6):1184–1189, June 2017. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7931733.

Balogh:2008:BDH

[BBHS08] J. Balogh, S. L. Bezrukov, L. H.

Harper, and A. Seress. On the bandwidth of 3-dimensional Hamming graphs. Theoretical Computer Science, 407(1–3):488–495, November 6, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bannai:2017:RDB

[BBZ17] Eiichi Bannai, Etsuko Bannai, and Yan Zhu. Relative t-designs in binary Hamming association scheme H(n, 2). Designs, Codes, and Cryptography, 84(1–2):23–53, July 2017. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic).

Bozorth:1953:MMS

[BH53] R. M. Bozorth and R. W. Hamming. Measurement of magnetostriction in single crystals. Physical Review, 89 (4):865–869, February 1953. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Bookstein:2001:FHD

[BKR01] Abraham Bookstein, Shmuel Tomi Klein, and Timo Raita. Fuzzy Hamming distance: A new dissimilarity measure (extended abstract). Lecture Notes in Computer Science, 2089:86–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/2089/20890086.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/2089/
20890086.pdf.

Bloomfield:1984:BRD

[Blo84] Peter Bloomfield. Book review: Digital Filters (2nd ed.) by R. W. Hamming. Journal of the American Statistical Association, 79(387):736–737, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.
org/stable/2288442.

Brainerd:1973:HMS

[Bra73] Barron Brainerd. The humanist meets statistics: [review of Introduction to Applied Numerical Analysis by R. W. Hamming]. Computers and the Humanities, 7(6):445–446, 1973. CODEN COHUAD. ISSN 0010-4817 (print), 1572-8412 (elec-

tronic). URL http://www.jstor.
org/stable/30199619.

Bultheel:1988:BRD

[Bul88] A. Bultheel. Book review: Digitale filter: Hamming R. W. Informationstechnologie VCH Verlagsgesellschaft, Weinheim, F. R. Germany, 1987, 275 pp., Price: DM 64, ISBN 3-527-26463-

9. Journal of Computational and Applied Mathematics, 24(3): N3, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/0377042788903068.

Bang:2008:SCH

[BvDK08] Sejeong Bang, Edwin R. van Dam, and Jack H. Koolen. Spectral characterization of the Hamming graphs. Linear Algebra and Its Applications, 429(11–12):2678– 2686, December 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chakrabarti:1999:LBC

[CCGL99] Amit Chakrabarti, Bernard Chazelle, Benjamin Gum, and Alexey Lvov. A lower bound on the complexity of approximate nearest-neighbor searching on the Hamming cube. In ACM [ACM99], pages 305–

311. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999. URL http:

//www.acm.org/pubs/articles/
proceedings/stoc/301250/p305-
chakrabarti/p305-chakrabarti.
pdf;   http://www.acm.org/pubs/
citations/proceedings/stoc/301250/

                                        p305-chakrabarti/. ACM order

number 508990.

Chang:2011:SFW

[CCLL11] Chin-Chen Chang, Kuo-Nan Chen, Chin-Feng Lee, and Li-Jen Liu. A secure fragile watermarking scheme based on chaos-and-Hamming code. The Journal of Systems and Software, 84(9):1462–1470, September 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0164121211000549.

Camion:1992:PCG

[CCM92] Paul Camion, Bernard Courteau, and Andre Montpetit. Un problème combinatoire dans les graphes de Hamming et sa solution en Scratchpad. (English: A combinatorial problem in Hamming graphs and its solution in Scratchpad). Rapports de recherche 1586, Institut National de Recherche en Informatique et en Automatique, Le Chesnay, France, January 1992. 12 pp.

Cormode:2002:CDS

[CDIM02] Graham Cormode, Mayur Datar, Piotr Indyk, and S. Muthukrishnan. Comparing data streams using Hamming norms (how to zero in). In Bernstein et al. [B^{+}02], pages 335–

345. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S10P02.pdf.

Ceruzzi:1980:NNO

[Cer80] Paul Ceruzzi. News and notices: Opening of Digital Computer Museum; Strachey Collection completed; anniversary of Hamming codes; CBI chooses permanent site; honorary degree awarded to Konrad Zuse. Annals of the History of Computing, 2(3):273–274, July/September 1980. CODEN AHCOE5. ISSN 0164-1239. URL http://dlib.

computer.org/an/books/an1980/
pdf/a3273.pdf;   http://www.
computer.org/annals/an1980/a3273abs.
htm.

Chung:1991:NGH

[Chu91] H. Chung. The 2-nd generalized Hamming weight of double-error correcting binary BCH codes and their dual codes. Lecture Notes in Computer Science, 539:118–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chung:2001:NAH

[Chu01] Sei-Jong Chung. Network architecture: Hamming codes and cyclic redundancy for transmission error correction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):48–50, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.
utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/
2001.bib.

Cheng:1995:UMH

[CM95] Jung-Fu Cheng and R. J. McEliece.

Unit-memory Hamming turbo codes. In Proceedings of 1995 IEEE International Symposium on Information Theory. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995.

Crescenzi:2002:HDC

[CR02] P. Crescenzi and G. Rossi. On the Hamming distance of constraint satisfaction problems. Theoretical Computer Science, 288(1):85–100, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chakrabarti:2011:OLB

[CR11] Amit Chakrabarti and Oded Regev. An optimal lower bound on the communication complexity of gap-Hamming-distance. In ACM [ACM11], pages 51–60. ISBN 1-4503-0691-8. LCCN ???? URL

                                        http://www.gbv.de/dms/tib-ub-
                                        hannover/63314455x..

Chakrabarti:2012:OLB

[CR12] Amit Chakrabarti and Oded Regev. An optimal lower bound on the communication complexity of gap-Hamming-distance. SIAM Journal on Computing, 41(5):1299–1317, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chang:2007:NIH

[CTL07] Chin-Chen Chang, Wei-Liang Tai, and Chia-Chen Lin. A novel image hiding scheme based on VQ and Hamming distance. Fundamenta Informaticae, 77(3):217–228, July 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cunningham:1973:RNM

[Cun73] W. J. Cunningham. Review: Numerical Methods for Scientists and Engineers (2nd ed.) by R. W. Hamming. American Scientist, 61(5): 605, ???? 1973.

Cusick:2013:FHW

[Cus13] Thomas W. Cusick. Finding Hamming weights without looking at truth tables. Cryptography and Communications: Discrete Structures, Boolean Functions and Sequences, 5(1):7–18, March 2013. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http://

link.springer.com/accesspage/
article/10.1007/s12095-012-0072-
8.

Camarero:2015:TCH

[CVB15] Cristóbal Camarero, Enrique Vallejo, and Ramón Beivide. Topological characterization of Hamming and dragonfly networks and its implications on routing. ACM Transactions on Architecture and Code Optimization, 11(4):39:1–39:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Davis:1963:RPN

[Dav63] Philip J. Davis. Recent publications: Numerical Methods for Scientists and Engineers, by R. W. Hamming. American Mathematical Monthly, 70(2):229, February 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Denisiuk:2017:LDE

[DG17] Aleksander Denisiuk and Michał Grabowski. Low distortion embedding of the Hamming space into a sphere with quadrance metric and k-means clustering of nominalcontinuous data. Fundamenta Informaticae, 153(3):221–233, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dijkstra:1976:DP

[Dij76a] Edsger W. Dijkstra. A discipline of programming. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976. ISBN 0-13-215871-X. xvii + 217 pp. LCCN QA76.6 .D54. With a foreword by C. A. R. Hoare.

Dijkstra:1976:EARb

[Dij76b] Edsger W. Dijkstra. An exercise attributed to R. W. Hamming. In A discipline of programming [Dij76a], chapter 17, pages 129–134. ISBN 0-13-215871-X. LCCN QA76.6 .D54. With a foreword by C. A. R. Hoare.

Das:1995:OEH

[DM95] S. K. Das and A. Mao. Optimal embeddings in the Hamming cube networks. Lecture Notes in Computer Science, 966:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denning:1997:BCN

[DM97] Peter J. Denning and Robert M. Metcalfe, editors. Beyond Calculation: The Next Fifty Years of Computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94932-1, 0-387-98588-3 (paperback), 1-4612-0685-5. xviii + 313 pp. LCCN QA76 .D348 1997.

Dougherty:2004:MDC

[DS04] Steven T. Dougherty and Keisuke Shiromoto. Maximum distance codes in Mat_{n,s}(Z_{k}) with a non-Hamming metric and uniform distributions. Designs, Codes, and Cryptography, 33(1):45–61, August 2004. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (elec-

tronic). URL http://ipsapp008.
kluweronline.com/IPS/content/
ext/x/J/4630/I/63/A/4/abstract.
htm.

Duren:1988:CMA

[Dur88] Peter Duren. A century of mathematics in America, volume 1 of History of mathematics. American Mathematical Society, Providence, RI, USA, 1988. ISBN 0-8218-0124-4. viii + 477 pp. LCCN QA27.U5 C46 1988.

Eggleton:1996:PSP

[EDD^{+}96] Roger B. Eggleton, Emeric Deutsch, Dragomir Z. Dokovic, Richard Hamming, Roger Pinkham, Fu-Chuen Chang, David Callan, and Hauke Reddmann. Problems and solutions: Problems: 10501–10507. American Mathematical Monthly, 103(2): 171–172, ???? 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Epstein:1972:NQF

[EH72] M. P. Epstein and R. W. Hamming. Noninterpolatory quadrature formulas. SIAM Journal on Numerical Analysis, 9(3):464–475, September 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ehrenborg:2021:FGH

[Ehr21] Richard Ehrenborg. Fano, Galois, Hamming and a card trick. College Mathematics Journal, 52 (4):274–280, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

                                        tandfonline.com/doi/full/10.
                                        1080/07468342.2021.1942678.

Ericson:1991:GCH

[EIZ91] T. Ericson, I. Ingemarsson, and V. Zinoviev. Group codes in Hamming spaces. In Proceedings. 1991 IEEE International Symposium on Information Theory, page 351. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991.

Ernvall:1985:NCH

[EKP85] Jarmo Ernvall, Jyrki Katajainen, and Martti Penttonen. NPcompleteness of the Hamming salesman problem. BIT (Nordisk tidskrift for informationsbehandling), 25(1):289–292, March 1985. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/openurl.asp?
                                        genre=article&issn=0006-3835&
                                        volume=25&issue=1&spage=289.

Ericson:1993:SCH

[EL93] T. Ericson and V. Levenshtein. Superimposed codes in Hamming space. In Proceedings. IEEE International Symposium on Information Theory, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993.

Elias:1955:CNC

[Eli55] Peter Elias. Coding for noisy channels. IRE Convention Record, 3(4):37–46, ???? 1955. URL

http://web.mit.edu/6.441/www/
reading/hd2.pdf.

Elias:1956:CTN

[Eli56] Peter Elias. Coding for two noisy channels. In Information Theory: Third London Symposium, pages 61–74. Butterworth Scientific, London, UK, 1956.

Elias:1974:CNC

[Eli74] Peter Elias. Coding for noisy channels. In Elwyn R. Berlekamp, editor, Key Papers in the Development of Coding Theory, IEEE Press selected reprint series. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1974. ISBN 0-87942-031-6, 0-87942-032-4 (paperback). LCCN QA268 .B47.

El-Qawasmeh:2000:PIH

[EQH00] Eyas El-Qawasmeh and Ismail Hmeidi. Performance investigation of Hamming Distance Bit Vertical Counter applied to access methods in information retrieval. Journal of the American Society for Information Science, 51(5):427–431, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fenwick:2015:NVL

[Fen15] P. Fenwick. A note on variablelength codes with constant Hamming weights. J.UCS: Journal of Universal Computer Science, 21(9): 1136–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://www.

                                        jucs.org/jucs_21_9/a_note_on_
                                        variable.

Fisher:1998:RHD

[Fis98] Lawrence M. Fisher. Richard Hamming, 82, dies; pioneer in digital technology. New York Times, ??(??):29, January 11, 1998. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

                                        URL https://search.proquest.
                                        com/hnpnewyorktimes/docview/
                                        109909823/.

Federico:2016:CTR

[FvdHH16] Lorenzo Federico, Remco van der Hofstad, and Tim Hulshof. Connectivity threshold for random subgraphs of the Hamming graph. Electronic Communications in Probability, 21(??):27:1–27:8, ???? 2016. CODEN ???? ISSN 1083-589X.

                                        URL https://projecteuclid.
                                        org/euclid.ecp/1457978024.

Gallager:2009:PEN

[Gal09] Robert G. Gallager. Peter Elias (November 26, 1923–December 7, 2001). Biographical memoirs — National Academy of Sciences of the United States of America, 91(??): 108–122, 2009. URL https://www.

                                        nap.edu/read/12776/chapter/7.

Gentle:1983:CSS

[Gen83] James E. Gentle, editor. Computer science and statistics: proceedings of the Fifteenth Symposium on the Interface, Houston, Texas, March 1983. North-Holland, Amsterdam, The Netherlands, 1983. ISBN 0-444-86688-4 (paperback). LCCN QA276.4 .S95 1983.

Grabowski:2008:BPS

[GF08] Szymon Grabowski and Kimmo Fredriksson. Bit-parallel string matching under Hamming distance in O(n⌈m/w⌉) worst case time. Information Processing Letters, 105 (5):182–187, February 29, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gravner:2015:BPH

[GHPS15] Janko Gravner, Christopher Hoffman, James Pfeiffer, and David Sivakoff. Bootstrap percolation on the Hamming torus. Annals of Applied Probability, 25(1): 287–323, February 2015. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/
euclid.aoap/1418740187.

Gilbert:1952:CSA

[Gil52] E. N. Gilbert. A comparison of signalling alphabets. The Bell System Technical Journal, 31 (3):504–522, May 1952. CO-DEN BSTJAN. ISSN 0005-8580.

URL http://bstj.bell-labs.
com/BSTJ/images/Vol31/bstj31-
3-504.pdf;   http://www.alcatel-
lucent.com/bstj/vol31-1952/articles/
bstj31-3-504.pdf.

Givens:1962:RSS

[Giv62] Wallace Givens. Reports of sections and societies: Mathematics (section a). Science, 135 (3503):547–549, February 16, 1962. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

                                        tronic). URL http://science.
                                        sciencemag.org/content/135/3503/
                                        547.full.pdf.

Groult:2002:ETR

[GLM02] Richard Groult, Martine Léonard, and Laurent Mouchard. Evolutive tandem repeats using Hamming distance. Lecture Notes in Computer Science, 2420:292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://link.springer.de/
                                        link/service/series/0558/bibs/
                                        2420/24200292.htm;   http://
                                        link.springer.de/link/service/
                                        series/0558/papers/2420/24200292.
                                        pdf.

Godsil:2015:HS

[GM15] Christopher Godsil and Karen Meagher. The Hamming scheme. In Erdős–Ko–Rado Theorems: Algebraic Approaches, pages 184–209. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-12844-7 (hardcover), 1-316-41495-7 (e-book). LCCN QA164 .G63 2016.

Guneri:2006:IGH

[GÖ06] Cem Güneri and Ferruh Özbudak. Improvements on generalized Hamming weights of some trace codes. Designs, Codes, and Cryptography, 39(2):215–231, May 2006.

CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0925-1022&
volume=39&issue=2&spage=215.

Golay:1949:NDC

[Gol49] Marcel J. E. Golay. Notes on digital coding. Proceedings of the Institute of Radio Engineers, 37 (??):657, June 1949. CODEN PIREAE. ISSN 0096-8390 (print), 2162-6634 (electronic). URL http:/

/www.maths.manchester.ac.uk/~
ybazlov/code/golay_paper.pdf;
 https://web.archive.org/web/
20161007122006/http://www.maths.
manchester.ac.uk/~ybazlov/code/
golay_paper.pdf. This short work

on error detecting codes predates that of [Ham50a]. It is cited in [War13, Chapter 15, pages 331ff], but is not mentioned in the 2003 first edition of that book.

Gottesman:1996:CQE

[Got96] Daniel Gottesman. Class of quantum error-correcting codes saturating the quantum Hamming bound. Physical Review A (Atomic, Molecular, and Optical Physics), 54:1862– 1868, September 1996. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

URL https://link.aps.org/doi/
10.1103/PhysRevA.54.1862.

Gillespie:2013:NTC

[GP13] Neil I. Gillespie and Cheryl E.

Praeger. Neighbour transitivity on codes in Hamming graphs. Designs, Codes, and Cryptography, 67 (3):385–393, June 2013. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1007/s10623-012-9614-5.

Grolmusz:2006:PCP

[Gro06] Vince Grolmusz. Pairs of codes with prescribed Hamming distances and coincidences. Designs, Codes, and Cryptography, 41(1):87–99, October 2006. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?
                                        genre=article&issn=0925-1022&
                                        volume=41&issue=1&spage=87.

Guizzo:2004:CPC

[Gui04] Erico Guizzo. Closing in on the perfect code [turbo codes]. IEEE Spectrum, 41(3):36–42, March 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (elec-

                                        tronic). URL https://spectrum.
                                        ieee.org/computing/software/
                                        closing-in-on-the-perfect-code.

Guttorp:1992:BRA

[Gut92a] Peter Guttorp. Book review: The Art of Probability for Scientists and Engineers by Richard W. Hamming. Technometrics, 34(2):242, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

                                        jstor.org/stable/1269264.

Guttorp:1992:BRB

[Gut92b] Peter Guttorp. Book review: The Art of Probability for Scientists and Engineers by Richard W. Hamming. Technometrics, 34(2):242, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269264.

Hamming:1942:SPBb

[Ham42a] Richard W. Hamming. Some Problems in the Boundary Value Theory of Linear Differential Equations. PhD thesis, University of Illinois at Urbana-Champaign, Urbana-Champaign, IL, USA, 1942.

26 pp. URL http://gateway.
proquest.com/openurl?url_ver=

Z39.88-2004. Thesis (Ph.D.)– University of Illinois at Urbana-Champaign.

Hamming:1942:SPBa

[Ham42b] Richard Wesley Hamming. Some problems in the boundary value theory of linear differential equations, 1942. Abstract of a thesis.

Hamming:1945:CMS

[Ham45a] R. W. Hamming. Convergent monotone series. American Mathematical Monthly, 52(2):70–72, February 1945. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamming:1945:DNT

[Ham45b] R. W. Hamming. Discussions and notes: The transcendental character of cos X. American Mathematical Monthly, 52(6):336–337, ???? 1945. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamming:1946:MS

[Ham46] R. W. Hamming. Monotone series.

American Journal of Mathematics, 68(??):133–136, ???? 1946. CODEN AJMAAN. ISSN 0002-9327 (print), 1080-6377 (electronic).

Hamming:1947:MNS

[Ham47] R. W. Hamming. Mathematical notes: Subseries of a monotone divergent series. American Mathematical Monthly, 54(8):462–463, October 1947. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See corrections [Ham48].

Hamming:1948:CSM

[Ham48] R. W. Hamming. Mathematical notes: Corrections: Subseries of monotone divergent series. American Mathematical Monthly, 55(6):360, June 1948. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [Ham47].

Hamming:1950:EDE

[Ham50a] R. W. Hamming. Error detecting and error correcting codes. The Bell System Technical Journal, 29(2):147–160, April 1950. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol29/bstj29-
2-147.pdf;   http://www.alcatel-
lucent.com/bstj/vol29-1950/articles/
bstj29-2-147.pdf.

Hamming:1950:OAC

[Ham50b] R. W. Hamming. Other aids to computation. Mathematics of Computation, 4(32):239–244, 1950. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). cited By 0.

Hamming:1951:RPG

[Ham51a] R. W. Hamming. Recent publications: Giant brains, or machines that think. American Mathematical Monthly, 58(4):276, ???? 1951. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamming:1951:SMS

[Ham51b] Richard W. Hamming. Some methods of solving hyperbolic and parabolic partial differential equations. In Proceedings, Computation Seminar, December 1949, pages 14– 23. International Business Machines Corporation, New York, NY, USA, 1951.

Hamming:1952:OAC

[Ham52] Richard W. Hamming. Other aids to computation. Mathematics of Computation, 6(40):248, October 1952. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hamming:1956:BRa

[Ham56a] R. W. Hamming. Book reviews.

Journal of the ACM, 3(2):112–113, April 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1956:BRb

[Ham56b] R. W. Hamming. Book reviews.

Journal of the ACM, 3(3):239, July 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1956:BRc

[Ham56c] R. W. Hamming. Book reviews.

Journal of the ACM, 3(4):376–378, October 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1957:BRa

[Ham57a] R. W. Hamming. Book reviews. Journal of the ACM, 4(1):93–94, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1957:BRb

[Ham57b] R. W. Hamming. Book reviews. Journal of the ACM, 4(2):219–220, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1957:BRc

[Ham57c] R. W. Hamming. Book reviews. Journal of the ACM, 4(3):362–366, July 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1957:BRd

[Ham57d] R. W. Hamming. Book reviews. Journal of the ACM, 4(4):530–533, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1957:CDC

[Ham57e] R. W. Hamming. Controlling the digital computer. The Scientific Monthly, 85(4):169–175, October 1957. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (elec-

                                        tronic). URL https://www.jstor.
                                        org/stable/21639.

Hamming:1958:BRa

[Ham58a] R. W. Hamming. Book reviews.

Journal of the ACM, 5(1):116, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1958:BRb

[Ham58b] R. W. Hamming. Book reviews.

Journal of the ACM, 5(2):197–203, April 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1958:BRc

[Ham58c] R. W. Hamming. Book reviews.

Journal of the ACM, 5(3):298–308, July 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1958:BRd

[Ham58d] R. W. Hamming. Book reviews.

Journal of the ACM, 5(4):389–396, October 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1959:BRa

[Ham59a] R. W. Hamming. Book reviews.

Journal of the ACM, 6(1):115–120, January 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1959:BRb

[Ham59b] R. W. Hamming. Book reviews.

Journal of the ACM, 6(2):306–312, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1959:BRc

[Ham59c] R. W. Hamming. Book reviews. Journal of the ACM, 6(3):443–458, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1959:BRd

[Ham59d] R. W. Hamming. Book reviews. Journal of the ACM, 6(4):548–556, October 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1959:SPC

[Ham59e] R. W. Hamming. Stable predictor– corrector methods for ordinary differential equations. Journal of the ACM, 6(1):37–47, January 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1959:SED

[Ham59f] Richard W. Hamming. Symposium on error detection and correction. In Information Processing, Proceedings of the International Conference on Information Processing, UNESCO, Paris 15– 20 June 1959, pages 487–491. UN-ESCO, Paris, France, 1959. URL

https://dblp.uni-trier.de/db/
conf/ifip/ifip1959.

Hamming:1960:BR

[Ham60] R. W. Hamming. Book reviews. Journal of the ACM, 7(1):80–86, January 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1961:MS

[Ham61a] R. Hamming. The mechanization of science. In Proceedings of the 1961 16th ACM National Meeting, ACM ’61, pages 31.101–31.104. ACM Press, New York, NY 10036, USA, 1961.

Hamming:1961:IIC

[Ham61b] R. W. Hamming. [intellectual implications of the computer revolution]. In ????, editor, [Man and the Computer: Proceedings of a Symposium at the the annual meeting of the American Association for the Advancement of Science held in Denver, CO, 27–29 December 1961], page ?? ????, ????, 1961.

Hamming:1962:CSS

[Ham62a] R. W. Hamming. Computers and simulation in systems engineering. IRE Transactions on Education, E-5(2):76–78, June 1962. ISSN 0893-7141.

Hamming:1962:NMS

[Ham62b] R. W. (Richard Wesley) Hamming.

Numerical Methods for Scientists and Engineers. International Series in Pure and Applied Mathematics. McGraw-Hill, New York, NY, USA, 1962. xvii + 411 pp. LCCN QA297 .H28.

Hamming:1963:IIC

[Ham63a] R. W. Hamming. Intellectual implications of the computer revolution. American Mathematical Monthly, 70(1):4–11, ???? 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamming:1963:SAS

[Ham63b] R. W. Hamming. State of the art in scientific computing. In Proceedings of the May 21–23, 1963, Spring Joint Computer Conference, AFIPS ’63 (Spring), pages 163–167. ACM Press, New York, NY 10036, USA,

1963. URL http://doi.acm.org/
10.1145/1461551.1461573.

Hamming:1964:BRE

[Ham64] R. W. Hamming. Book review: Edward L. Stiefel, An Introduction to Numerical Mathematics, Academic Press, Inc., New York, 1963, x + 286 p., 24 cm. Mathematics of Computation, 18(88):684, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.jstor.
org/stable/2002969.

Hamming:1965:BRE

[Ham65a] R. W. Hamming. Book review: Electronic Analog and Hybrid Computers. Granino A. Korn and Theresa M. Korn. McGraw-Hill, New York, 1964. xxiv + 584 pp. Illus. $17.50. Science, 148(3668):356– 361, April 16, 1965. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Hamming:1965:ENA

[Ham65b] R. W. Hamming. Book review: Elements of Numerical Analysis, by Peter Henrici. Mathematics of Computation, 19(91):516–517, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.jstor.
org/stable/2003710.

Hamming:1965:C

[Ham65c] R. W. Hamming. Contributors.

IEEE Transactions on Electronic Computers, EC-14(5):742, October 1965. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038569.

Hamming:1965:IC

[Ham65d] R. W. Hamming. Impact of computers. American Mathematical Monthly, 72(2, part II):1–7, ???? 1965. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamming:1965:LEM

[Ham65e] R. W. Hamming. Letters to the editor: Mathematics vs. numerical analysis. Science, 149(3681):245, July 16, 1965. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Hamming:1965:NAV

[Ham65f] R. W. Hamming. Numerical analysis vs. mathematics. Science, 148(3669):473–475, April 23, 1965. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://science.
sciencemag.org/content/148/3669/
473.

Hamming:1966:RRP

[Ham66a] R. Hamming. Review of Random-Process Simulation and Measurements (Korn, G. A.; 1966). IEEE Transactions on Information Theory, 12(4):489, October 1966. CO-DEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Hamming:1966:DJ

[Ham66b] R. W. Hamming. Dear John. Simulation, 6(6):vi, June 1966. CODEN SIMUA2. ISSN 0037-5497 (print), 1741-3133 (elec-

                                        tronic). URL http://journals.
                                        sagepub.com/doi/pdf/10.1177/
                                        003754976600600613.

Hamming:1966:SNA

[Ham66c] R. W. Hamming. Statistics in numerical analysis. Communications of the ACM, 9(7):482, July 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hamming:1966:CCR

[Ham66d] R. W. (Richard Wesley) Hamming. Calculus and the Computer Revolution. Committee on the Undergraduate Program in Mathematics. CUPM monograph. Committee on the Undergraduate Program in Mathematics, Berkeley, CA, USA, 1966. 24 pp. LCCN QA76.5 .H353.

Hamming:1968:BRC

[Ham68a] R. W. Hamming. Book review: Computation: Finite and Infinite Machines. Marvin L. Minsky. Prentice-Hall, Englewood Cliffs, N.J., 1967. xviii + 317 pp., illus. $12. Prentice-Hall Series in Automatic Computation. Science, 159 (3818):966–967, March 1, 1968. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Hamming:1968:BRF

[Ham68b] R. W. Hamming. Book review: Finding optima. Foundations of Optimization. Douglass J. Wilde and Charles S. Beightler. Prentice-Hall, Englewood Cliffs, N.J., 1967. xiv + 480 pp., illus. $12.50. Science, 159 (3821):1347, March 22, 1968. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Hamming:1968:CCR

[Ham68c] R. W. (Richard Wesley) Hamming.

Calculus and the Computer Revolution. Houghton Mifflin, Boston, 1968. x + 72 pp. LCCN QA76.5 .H353 1968.

Hamming:1968:BRT

[Ham68d] Richard W. Hamming. Book review: The Theory of Gambling and Statistical Logic, Richard A. Epstein. Academic Press, New York, 1967. xvi + 492 pp. Science, 161 (3844):878, August 30, 1968. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Hamming:1969:OMV

[Ham69a] R. W. Hamming. One man’s view of computer science. Journal of the ACM, 16(1):3–12, January 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1969:RCS

[Ham69b] Richard W. Hamming. Remote controlled system for reducing distortion. US Patent 3,440,539A., April 22, 1969. Patent filed 16 December 1964.

Hamming:1970:MCT

[Ham70a] R. Hamming. Modern control theory. IEEE Transactions on Automatic Control, 15(2):245, April 1970. CODEN IETAA9. ISSN 0018-9286 (print), 1558-2523 (electronic).

Hamming:1970:CNE

[Ham70b] R. W. Hamming. Classroom notes: An elementary discussion of the transcendental nature of the elementary transcendental functions. American Mathematical Monthly, 77(3):294–297, March 1970. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamming:1970:EDT

[Ham70c] R. W. Hamming. An elementary discussion of the transcendental nature of the elementary transcendental functions. American Mathematical Monthly, 77(3):294–297, March 1970. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamming:1970:DN

[Ham70d] R. W. Hamming. On the distribution of numbers. The Bell System Technical Journal, 49(8): 1609–1625, October 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol49/bstj49-
                                        8-1609.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/articles/
                                        bstj49-8-1609.pdf.

Hamming:1971:CIF

[Ham71a] R. W. Hamming. A class of integration formulas. American Mathematical Monthly, 78(5):518–522, May 1971. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamming:1971:MNC

[Ham71b] R. W. Hamming. Mathematical notes: a class of integration formula. American Mathematical Monthly, 78(5):518–522, May 1971. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamming:1971:CS

[Ham71c] R. W. (Richard Wesley) Hamming.

Computers and Society. McGraw-Hill computer science series. Mc-Graw-Hill, New York, NY, USA, 1971. ISBN 0-07-028593-7. x + 284 pp. LCCN QA76.5 .H354.

Hamming:1971:IAN

[Ham71d] R. W. (Richard Wesley) Hamming.

Introduction to Applied Numerical Analysis. McGraw-Hill computer science series. McGraw-Hill, New York, NY, USA, 1971. ISBN 0-07-025889-9. x + 331 pp. LCCN QA297 .H275.

Hamming:1972:CS

[Ham72] Richard W. Hamming. Computers and Society. McGraw-Hill, New York, NY, USA, 1972. ISBN 0-07-028593-7, 0-07-028593-4. x + 284 pp. LCCN QA76.5 .H354. McGraw-Hill Computer Science Series.

Hamming:1973:NMS

[Ham73a] R. W. Hamming. Numerical Methods for Scientists and Engineers. McGraw-Hill, New York, NY, USA, second edition, 1973. ISBN 0-07-025887-2. ix + 721 pp. International Series in Pure and Applied Mathematics.

Hamming:1973:RTS

[Ham73b] R. W. Hamming. Role of the technical societies in the field of computer measurement. Technical Report Special Publication 424, U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1973. 151– 153 pp.

Hamming:1973:BRN

[Ham73c] Richard W. Hamming. Book review: Numerical Analysis and Computation: Theory and Practice (E. K. Blum). SIAM Review, 15(4):807, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hamming:1974:FAN

[Ham74] R. W. Hamming. The frequency approach to numerical analysis. In Scaife [Sca74], pages 151–175. ISBN 0-12-621150-7. LCCN QA297

.S86.  URL http://catalog.
hathitrust.org/Record/000575557.

Published for the Royal Irish Academy.

Hamming:1975:HDY

[Ham75a] R. W. Hamming. How do you know the simulation is relevant? Simulation, 25(5):163–167, 1975. CODEN SIMUA2. ISSN 0037-5497 (print), 1741-3133 (elec-

tronic). URL http://journals.
sagepub.com/doi/pdf/10.1177/
003754977502500505.

Hamming:1975:PCS

[Ham75b] R. W. Hamming. A philosophy for computer science or my prejudices and confusions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7 (4):16–18, December 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamming:1977:DF

[Ham77] R. W. (Richard Wesley) Hamming.

Digital Filters. Prentice-Hall signal processing series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1977. ISBN 0-13-212571-4. xii + 226 pp. LCCN QA297 .H26.

Hamming:1979:RZW

[Ham79] Richard W. Hamming. Reviews: Z. W. Pylyshyn: Perspectives on the Computer Revolution; capsule reviews. Annals of the History of Computing, 1 (2):100, 208, October/December 1979. CODEN AHCOE5. ISSN 0164-1239. URL http://dlib.

computer.org/an/books/an1979/
pdf/a2100.pdf;   http://www.
computer.org/annals/an1979/a2100abs.
htm.

Hamming:1980:UEM

[Ham80a] R. W. Hamming. The unreasonable effectiveness of mathematics. American Mathematical Monthly, 87(2):81–90, February 1980. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.
org/stable/2321982. See further

analysis [Isl22].

Hamming:1980:WWK

[Ham80b] R. W. Hamming. We would know what they thought when they did it. In Metropolis et al. [MHR80], pages 3–9. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Hamming:1980:CIT

[Ham80c] Richard Wesley Hamming. Coding and Information Theory. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1980. ISBN 0-13-139139-9. xii + 239 pp. LCCN QA268 .H35 1980. US$19.95.

Hamming:1983:F

[Ham83a] R. W. Hamming. Future. In Gentle [Gen83], pages 3–6. ISBN 0-444-86688-4 (paperback). LCCN QA276.4 .S95 1983.

Hamming:1983:DF

[Ham83b] R. W. (Richard Wesley) Hamming. Digital Filters. Prentice-Hall signal processing series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1983. ISBN 0-13-212506-4. xiv + 257 pp. LCCN QA297 .H26 1983.

Hamming:1984:CDM

[Ham84] R. W. Hamming. Calculus and discrete mathematics . . .. College Mathematics Journal, 15(5): 388–389, November 1984. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

                                        URL http://www.tandfonline.
                                        com/doi/abs/10.1080/00494925.
                                        1984.11972813.

Hamming:1985:UEM

[Ham85a] R. W. Hamming. The unreasonable effectiveness of mathematics [Amer.

Math. Monthly 87 (1980), no. 2, 81–90; MR0559142 (81e:00015)]. In Mickens [Mic85], pages 15–29. ISBN 0-442-26077-6. LCCN QC20 .M36 1985.

Hamming:1985:MMA

[Ham85b] R. W. (Richard Wesley) Hamming.

Methods of Mathematics Applied to Calculus, Probability, and Statistics. Prentice-Hall series in computational mathematics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-578899-4. xviii + 857 pp. LCCN QA37.2 .H26 1985. US$40.00.

Hamming:1986:CIT

[Ham86a] R. W. (Richard Wesley) Hamming. Coding and Information Theory. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1986. ISBN 0-13-139072-4, 0-13-139139-9. xii + 259 pp. LCCN QA268 .H35 1986. US$36.95.

Hamming:1986:NMS

[Ham86b] R. W. (Richard Wesley) Hamming.

Numerical Methods for Scientists and Engineers. Dover, New York, NY, USA, second edition, 1986. ISBN 0-486-65241-6 (paperback). ix + 721 pp. LCCN QA297 .H28 1986.

US$14.95. URL http://www.loc.
gov/catdir/description/dover032/
86016226.html.

Hamming:1986:YYR

[Ham86c] Richard W. Hamming. You and your research. Web transcript and video of talk at Bellcore., March 7, 1986. URL http://www.

maultech.com/chrislott/misc/
kaiser.html;   http://www.paulgraham.

                                        com/hamming.html;   https://
                                        www.youtube.com/watch?v=8j3zEcusTuA.

Hamming:1987:OMV

[Ham87a] R. W. Hamming. One man’s view of computer science. In Ashenhurst [Ash87], pages 3–12. ISBN 0-201-07794-9. LCCN QA76.24 .A33 1987. ACM Turing Award lecture.

Hamming:1987:ICG

[Ham87b] R. W. (Richard Wesley) Hamming. Information und Codierung. (German) [Coding and Information Theory]. Informationstechnologie. VCH, Weinheim, Federal Republic of Germany, 1987. ISBN 3-527-26611-9. xii + 269 pp. LCCN QA268 H224.

Hamming:1987:DF

[Ham87c] Richard W. Hamming. Digitale Filter. Informationstechnologie. VCH Verlagsgesellschaft, Weinheim, West Germany, 1987. ISBN 3-527-26463-9. xii + 277 pp. Übers. aus dem Engl. von Bernhard H. Steinebrunner und Joachim Durzok) (Digital filters).

Hamming:1988:UM

[Ham88a] R. W. Hamming. The use of mathematics. In A century of mathematics in America [Dur88], pages 429–

437. ISBN 0-8218-0124-4. LCCN

QA27.U5 C46 1988.

Hamming:1988:RFC

[Ham88b] Richard Hamming. Reviews: A First Course in Coding Theory. American Mathematical Monthly, 95(8):786–787, ???? 1988. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamming:1988:RTL

[Ham88c] Richard W. Hamming. Reviews:

Toward a lean and lively calculus: Report of the Conference/Workshop to Develop Curriculum and Teaching Methods for Calculus at the College Level. American Mathematical Monthly, 95(5):466–471, ???? 1988. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamming:1989:TBP

[Ham89a] R. W. Hamming. The tennis ball paradox. Mathematics Magazine, 62 (4):268–269, ???? 1989. CODEN MAMGA8. ISSN 0025-570X. URL

http://www.jstor.org/stable/
2689769?origin=pubexport.

Hamming:1989:DF

[Ham89b] R. W. (Richard Wesley) Hamming.

Digital Filters. Prentice Hall signal processing series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1989. ISBN 0-13-212812-8, 0-13-212895-0. xiv + 284 pp. LCCN QA297 .H26 1989. US$35.25.

Hamming:1989:IAN

[Ham89c] R. W. (Richard Wesley) Hamming.

Introduction to Applied Numerical Analysis. Hemisphere Pub. Corp., New York, NY, USA, 1989. ISBN 0-89116-865-6. x + 331 pp. LCCN QA297 .H275 1989. US$70.00 (est.).

Hamming:1990:BRU

[Ham90a] R. W. Hamming. Book review:

[untitled]. Journal of the American Statistical Association, 85(410):592, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

                                        tronic). URL http://www.jstor.
                                        org/stable/2289807.

Hamming:1990:EDE

[Ham90b] R. W. Hamming. Error detecting and error correcting codes. In Computer Arithmetic [Swa90], pages 16– 29. ISBN 0-8186-8945-5. LCCN QA76.9 .C62C66 1990. Reprint of [Ham50a].

Hamming:1990:FH

[Ham90c] R. W. Hamming. Foreword by Hamming. In Digital Signal Analysis [SH90], pages ??–? ISBN 0-13-213117-X, 0-13-211772-X (paperback). LCCN TK5102.5 .S698 1990.

Hamming:1990:FS

[Ham90d] Richard W. Hamming. The future of statistics. The American Statistician, 44(2):133–135, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.jstor.
                                        org/stable/2684152.

Hamming:1991:HWL

[Ham91a] R. W. Hamming. How I was led to the frequency approach. Digital Signal Processing, 1(3):180–184, July 1991. CODEN DSPREJ. ISSN 1051-2004 (print), 1095-4333 (electronic).

Hamming:1991:APS

[Ham91b] R. W. (Richard Wesley) Hamming. The Art of Probability — for Scientists and Engineers. Addison-Wes-ley, Reading, MA, USA, 1991. ISBN 0-201-40686-1, 0-201-51058-8. xvi + 344 pp. LCCN QA273 .H3544 1991.

Hamming:1992:RJN

[Ham92a] R. W. Hamming. Review of John von Neumann and the Origins of Modern Computing (Aspray, W; 1990). Computer, 25(10):126–127, October 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamming:1992:RMF

[Ham92b] R. W. Hamming. Review of Mathematical Foundations of Computer Science, Volume 1 (Fejer, P. A. and Simovici D. A; 1991). Computer, 25 (1):134–135, January 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamming:1993:HDD

[Ham93] R. W. Hamming. Hamming distance and its discovery. The UMAP Journal: The Journal of Undergraduate Mathematics and its Applications, 14(4):??, ???? 1993. ISSN 0197-3622 (print), 1938-338X (elec-

tronic). URL http://www.comap.
com/product/?idx=1252.

Hamming:1994:SEB

[Ham94] R. W. Hamming. Science is in the eye of the beholder. Computers and Chemistry, 18(3):353–357, September 1994. CODEN COCHDK. ISSN 0097-8485 (print), 1879-0763 (electronic).

Hamming:1995:DFP

[Ham95a] R. W. Hamming. Digital filters, part 1. Web video of lecture., April 27, 1995. URL https://www.

                                        youtube.com/watch?v=2wO8ZHnn3Q4.

Hamming:1995:ECC

[Ham95b] R. W. Hamming. Error-correcting codes. Web video of lecture., April

                                        21, 1995. URL https://www.
                                        youtube.com/watch?v=BZh07Ew32UA.

Hamming:1995:IAD

[Ham95c] R. W. Hamming. Introduction to the art of doing science and engineering: Learning to learn. Web lecture video, March 28, 1995.

                                        URL https://www.youtube.com/
                                        watch?v=AD4b-52jtos.

Hamming:1996:FSE

[Ham96] R. W. Hamming. Foreword: Software engineering. Journal of Systems Integration, 6(1–2):5–7, March 1996.

Hamming:1997:HTA

[Ham97a] R. W. Hamming. How to think about trends. In Denning and Metcalfe [DM97], pages 65–74. ISBN 0-387-94932-1, 0-387-98588-3 (paperback), 1-4612-0685-5. LCCN QA76 .D348 1997.

Hamming:1997:ADS

[Ham97b] R. W. (Richard Wesley) Hamming. The Art of Doing Science and Engineering: Learning to Learn. Gordon and Breach, Amsterdam, The Netherlands, 1997. ISBN 90-5699-500-6, 90-5699-501-4 (paperback), 0-203-45071-X (e-book), 0-203-45913-X (Adobe eReader). vii + 218 pp. LCCN Q175 .H263 1997. URL http:/

                                        /site.ebrary.com/id/10166593;

http://www.tandfebooks.com/
isbn/9781135301033.

Hamming:1998:MDP

[Ham98a] R. W. Hamming. Mathematics on a distant planet. American Mathematical Monthly, 105(7):640–650, August/September 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamming:1998:DF

[Ham98b] R. W. (Richard Wesley) Hamming.

Digital filters. Dover, New York, NY, USA, third edition, 1998. ISBN 0-486-65088-X (paperback). xiv + 284 pp. LCCN QA297 .H26

1998. URL http://www.loc.gov/
catdir/description/dover032/
97051370.html;   http://www.
loc.gov/catdir/toc/dover031/
97051370.html.

Hamming:1999:VH

[Ham99] R. W. Hamming. Vorwort von Hamming. In Stearns and Hush [SH99], pages 13–14.

Hamming:2000:FPI

[Ham00] R. W. Hamming. The future of physics: information hiding. In Hecker and Rota [HR00], pages 73– 81. ISBN 1-4612-0777-0 (e-book), 1-4612-6898-2 (print), 0-8176-3856-3, 3-7643-3856-3. LCCN Q160.2 .E77 2000.

Hamming:2001:MDP

[Ham01] R. W. Hamming. Mathematics on a distant planet. Pokroky Matematiky, Fyziky a Astronomie, 46(3): 219–231, 2001. CODEN PMFAA4. ISSN 0032-2423.

Hamming:2004:MMA

[Ham04] R. W. (Richard Wesley) Hamming. Methods of Mathematics Applied to Calculus, Probability, and Statistics. Dover books on mathematics. Dover, New York, NY, USA, 2004. ISBN 0-486-43945-3 (paperback). xviii + 857 pp. LCCN QA37.3 .H35 2004. URL http://www.

loc.gov/catdir/enhancements/
fy0618/2004053443-d.html.

Hamming:2009:FCG

[Ham09] Richard Hamming. Fehlerkorrigierende Codes. (German) [errorcorrecting codes]. In Historische Notizen zur Informatik. (German) [Historical notes on computer science] [Bau09], pages 337–344. ISBN 3-540-85790-7. LCCN QA76.17 .B38 2009.

Hamming:2012:IAN

[Ham12a] R. W. (Richard Wesley) Hamming. Introduction to Applied Numerical Analysis. Dover books on mathematics. Dover, New York, NY, USA, 2012. ISBN 0-486-48590-0 (paperback). x + 331 pp. LCCN QA297 .H275 2012.

Hamming:2012:CEC

[Ham12b] Richard Hamming. Classical errorcorrecting codes: Machines should work. people should think. In Marinescu and Marinescu [MM12], pages 345–454. ISBN 0-12-383874-6 (hardcover). LCCN QA76.889 .M363 2012.

Hamming:2015:EDE

[Ham15a] R. W. Hamming. Error detecting and error correcting codes.

In Swartzlander, Jr. [Swa15b], pages 3–16. ISBN 981-4641-46-4 (hardcover), 981-4641-47-2 (ebook). LCCN QA76.6 .C633 2015 vol. 2.

Hamming:2015:DN

[Ham15b] R. W. Hamming. On the distribution of numbers. In Swartzlander, Jr. [Swa15a], pages 321–337. ISBN 981-4651-56-7 (vol. 1; hardcover), 981-4651-57-5, 981-4641-47-2 (e-book). LCCN QA76.6 .C633 2015 vol. 1.

Hamming:2021:EDE

[Ham21] R. W. Hamming. Error detecting and error correcting codes (1950). In Ideas That Created the Future: Classic Papers of Computer Science [Lew21], chapter 13, pages 135–146. ISBN 0-262-04530-3. LCCN Q124.6-127.2.

Harper:2003:BHG

[Har03] L. H. Harper. On the bandwidth of a Hamming graph. Theoretical Computer Science, 301(1–3): 491–498, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hemming:1968:CCM

[Hem68] R. V. Hemming. Chislennye metody dlya nauchnykh rabotnikov i inzhenerov (Russian) [Numerical methods for scientists and engineers]. Translated from the English by V. L. Arlazarov, G. S. Razina and A. V. Uskov. Edited by R. S. Guter. Izdat. “Nauka”, Moscow, USSR, 1968. 400 pp.

Hamming:1951:EDC

[HH51] Richard W. Hamming and Bernard D. Holbrook. Error-detecting and correcting system. US Patent 2,552,629A., May 15, 1951. URL

                                        https://patents.google.com/patent/
                                        US2552629A/en. Patent filed 11

January 1950.

Hopkins:1957:CR

[HH57] I. L. Hopkins and R. W. Hamming. On creep and relaxation. Journal of Applied Physics, 28(8):906–909, August 1957. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. See note [HH58].

Hopkins:1958:NPC

[HH58] I. L. Hopkins and R. W. Hamming. Note on paper On Creep and Relaxation. Journal of Applied Physics, 29(4):742, April 1958. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. See [HH57].

Heidenreich:1965:NEE

[HH65] R. D. Heidenreich and R. W. Hamming. Numerical evaluation of electron image phase contrast. The Bell System Technical Journal, 44(2):207–233, February 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol44/bstj44-
                                        2-207.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/articles/
                                        bstj44-2-207.pdf.

Hirst:2001:SSI

[HH01] Simon Hirst and Bahram Honary. A simple soft-input/soft-output decoder for Hamming codes. Lecture Notes in Computer Science, 2260:38–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/2260/22600038.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/2260/
22600038.pdf.

Hamming:1978:MFU

[HK78] Richard W. Hamming and James F.

Kaiser. Multisection filter using inflected amplitude change function to sharpen its band-edge responses. US Patent 4,074,212., February 14,

1978. URL https://www.google.
com/patents/US4074212. Patent

filed 16 December 1976.

Homma:2009:SGH

[HK09] Masaaki Homma and Seon Jeong Kim. The second generalized Hamming weight for two-point codes on a Hermitian curve. Designs, Codes, and Cryptography, 50(1):1–40, January 2009. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://www.

springerlink.com/openurl.asp?
genre=article&issn=0925-1022&
volume=50&issue=1&spage=1.

Honkala:2002:CIP

[HL02] Iiro Honkala and Antoine Lobstein.

On the complexity of the identification problem in Hamming spaces. Acta Informatica, 38(11/12):839– 845, November 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Honkala:2001:CIS

[HLR01] Iiro Honkala, Tero Laihonen, and Sanna Ranto. On codes identifying sets of vertices in Hamming spaces. Designs, Codes, and Cryptography, 24(2):193–204, October 2001. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (elec-

                                        tronic). URL http://www.wkap.
                                        nl/oasis.htm/353906.

Hamming:1965:NLB

[HM65] R. W. Hamming and W. L. Mammel. A note on the location of the binary point in a computing machine. IEEE Transactions on Electronic Computers, EC-14(2):260–261, April 1965. CODEN IEECA8. ISSN 0367-

                                        7508. URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?tp=&
                                        arnumber=4038414.

Houghton:1997:EEC

[Hou97a] A. D. Houghton. The Engineer’s Error Coding Handbook. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 1-4613-0447-4. 272 pp. LCCN TK7888.4.

Houghton:1997:HC

[Hou97b] A. D. Houghton. Hamming codes. In The Engineer’s Error Coding Handbook [Hou97a], chapter 4, pages 27–34. ISBN 1-4613-0447-4. LCCN TK7888.4.

Hamming:1966:CIF

[HP66] Richard W. Hamming and Roger S. Pinkham. A class of integration formulas. Journal of the ACM, 13 (3):430–438, July 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Han:2015:NLB

[HPL15] Hongyu Han, Daiyuan Peng, and Xing Liu. New lower bounds on the aperiodic Hamming correlations of frequency hopping sequences with low hit zone. Designs, Codes, and Cryptography, 75(1):157–174, April 2015. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://link.

springer.com/article/10.1007/
s10623-013-9900-x.

Han:2017:CLH

[HPP^{+}17] Hongyu Han, Daiyuan Peng, Udaya Parampalli, Zheng Ma, and Hongbin Liang. Construction of low-hit-zone frequency hopping sequences with optimal partial Hamming correlation by interleaving techniques. Designs, Codes, and Cryptography, 84(3):401–414, September 2017. CODEN DC-CREC. ISSN 0925-1022 (print), 1573-7586 (electronic).

He:2004:GCH

[HPR04] Matthew X. He, Sergei V. Petoukhov, and Paolo E. Ricci. Genetic code, Hamming distance and stochastic matrices. Bulletin of Mathematical Biology, 66(5):1405–1421, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.
1016/j.bulm.2004.01.002.pdf.

Hamming:1998:PSS

[HPS98] Richard Hamming, Roger Pinkham, and Richard Stong. Problems and solutions: Solutions: Another Polya urn scheme: 10504. American Mathematical Monthly, 105(2):181– 182, ???? 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hecker:2000:EFH

[HR00] Siegfried S. Hecker and Gian-Carlo Rota, editors. Essays on the Future: in honor of Nick Metropolis. Birkhäuser Boston Inc., Cambridge, MA, USA, 2000. ISBN 1-4612-0777-0 (e-book), 1-4612-6898-2 (print), 0-8176-3856-3, 3-7643-3856-3. xvii + 276 pp. LCCN Q160.2 .E77 2000.

Hamming:1954:RCS

[HRSC54] R. W. Hamming, Sibyl Rock, A. L. Samuel, and W. C. Carter. Redundancy checking for small digital computers. In Proceedings of the December 8–10, 1954, Eastern Joint Computer Conference: Design and Application of Small Digital Computers, AIEE-IRE ’54 (Eastern), pages 56–57. ACM Press, New York, NY 10036, USA, 1954. URL http:/

/doi.acm.org/10.1145/1455270.
1455283.

Hamming:1955:RCS

[HRSC55] R. W. Hamming, Sibyl Rock, A. L. Samuel, and W. C. Carter. Redundancy checking for small digital computers. In Proceedings of the Eastern Joint Computer Conference, Philadelphia, December 8–10, 1954, pages 56–57. AFIPS Press, Montvale, NJ, USA, 1955.

Hamming:1979:DF

[HS79] R. W. Hamming and S. D. Stearns.

Digital filters. IEEE Transactions on Systems, Man, and Cybernetics, 9(1):67, January 1979. CODEN ISYMAW. ISSN 0018-9472 (print), 2168-2909 (electronic).

Horimoto:2001:GHW

[HS01] H. Horimoto and K. Shiromoto.

On generalized Hamming weights for codes over finite chain rings. Lecture Notes in Computer Science, 2227:141–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/2227/22270141.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/2227/
22270141.pdf.

Huang:2006:CCH

[HSZZ06] Wei Huang, Yaoyun Shi, Shengyu Zhang, and Yufan Zhu. The communication complexity of the Hamming distance problem. Information Processing Letters, 99(4):149– 153, August 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Houselander:1990:IHB

[HT90] P. Houselander and J. T. Taylor.

Improving the Hamming binary associative memory. Electronics Letters, 26(11):705–707, May 1990. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Huning:1993:NSA

[Hun93] H. Huning. A node splitting algorithm that reduces the number of connections in a Hamming distance classifying network. Lecture Notes in Computer Science, 686:102–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamming:1980:DF

[HY80] R. W. Hamming and C. K. Yuen. Digital filters. IEEE Transactions on Systems, Man, and Cybernetics, 10(10):689, October 1980. CODEN ISYMAW. ISSN 0018-9472 (print), 2168-2909 (electronic).

Heng:2017:EHW

[HY17] Ziling Heng and Qin Yue. Evaluation of the Hamming weights of a class of linear codes based on Gauss sums. Designs, Codes, and Cryptography, 83(2):307–326, May 2017. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://

link.springer.com/accesspage/
article/10.1007/s10623-016-0222-
7;   http://link.springer.com/
article/10.1007/s10623-016-0222-
7.

IEEE:2018:IRW

[IEE18] IEEE. IEEE Richard W. Hamming Medal. IEEE Web site,

2018. URL https://www.ieee.
org/about/awards/medals/hamming.
html;   https://www.ieee.org/
content/dam/ieee-org/ieee/web/
org/about/hamming_rl.pdf.

Imrich:1997:RHG

[IK97] Wilfried Imrich and Sandi Klavžar.

Recognizing Hamming graphs in linear time and space. Information Processing Letters, 63(2):91–95, August 13, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Itoh:1986:RSC

[IS86] Yoshiaki Itoh and Herbert Solomon.

Random sequential coding by Hamming distance. Journal of Applied Probability, 23(3):688–695, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.
org/stable/3214007.

Irwin:1987:PSC

[IS87] Mary Jane Irwin and Renato Stefanelli, editors. Proceedings: 8th Symposium on Computer Arithmetic, May 19–21, 1987, Villa Olmo, Como, Italy. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 1987. ISBN 0-8186-0774-2 (paperback), 0-8186-4774-4 (microfiche), 0-8186-8774-6 (case). LCCN QA 76.9 C62 S95 1987.

Isaka:2005:HRS

[Isa05] M. Isaka. High-rate serially concatenated codes using Hamming codes. In IEEE International Conference on Communications, 2005. ICC 2005. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

Islami:2022:UEM

[Isl22] Arezoo Islami. The unreasonable effectiveness of mathematics: From Hamming to Wigner and back again. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://

link.springer.com/article/10.
1007/s10701-022-00592-8. See

[Ham80a].

Izadikhah:2009:UHD

[Iza09] Mohammad Izadikhah. Using the Hamming distance to extend TOP-SIS in a fuzzy environment. Journal of Computational and Applied Mathematics, 231(1):200–207, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/S0377042709000533.

Jain:2007:CTB

[Jai07] Sapna Jain. Campopiano-type bounds in non-Hamming array coding. Linear Algebra and Its Applications, 420(1):135–159, January 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jarrous:2013:SCF

[JP13] Ayman Jarrous and Benny Pinkas. Secure computation of functionalities based on Hamming distance and its application to computing document similarity. International Journal of Applied Cryptography. IJACT, 3(1):21–46, 2013. CODEN ???? ISSN 1753-0563 (print), 1753-0571 (electronic).

Juncosa:1962:NMS

[Jun62] M. L. Juncosa. Book review:

Applied mathematics: Numerical Methods for Scientists and Engineers. Richard W. Hamming. McGraw-Hill, New York, 1962. 411 pp. Illus. $11. Science, 138(3545): 1091, December 7, 1962. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Kaiser:2009:SRL

[Kai09] J. F. Kaiser. Richard Hamming — you and your research. In Aslak Tveito, A. M. (Are Magnus) Bruaset, and Olav Lysne, editors, Simula Research Laboratory: by thinking constantly about it, pages 37– 60. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., October 2009. ISBN 3-642-01155-1, 3-642-01156-X. LCCN T177.N7 S56 2009.

Karon:1977:RIA

[Kar77] John Karon. Reviews: Introduction to Applied Numerical Analysis, by Richard W. Hamming; Numerical Methods with FORTRAN IV Case Studies, by William S. Dorn and Daniel D. McCracken; Numerical Quadrature and Solution of Ordinary Differential Equations, by A. H. Stroud. American Mathematical Monthly, 84(4):304–307, April 1977. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kelly:2018:AIM

[Kel18a] Cindy Kelly. AtomicHeritage interview with Martin Mandelberg. YouTube 52m video., May 7, 2018. URL https:

                                        //www.manhattanprojectvoices.
                                        org/oral-histories/martin-mandelbergs-
                                        interview;   https://www.youtube.
                                        com/watch?v=cXLI-UpEAJo.

Kelly:2018:MMI

[Kel18b] Cindy Kelly. Martin Mandelberg’s interview. Web document., March 16, 2018. URL https:

                                        //www.manhattanprojectvoices.
                                        org/oral-histories/martin-mandelbergs-
                                        interview.

Kaiser:1977:SRS

[KH77] J. Kaiser and R. Hamming. Sharpening the response of a symmetric nonrecursive filter by multiple use of the same filter. In ICASSP ’77. IEEE International Conference on Acoustics, Speech, and Signal Processing, volume 2, pages 82–

85. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 1977.

Khemming:1972:CCM

[Khe72] R. V. Khemming. Chislennye metody dlya nauchnykh rabotnikov i inzhenerov (Russian) [Numerical methods for scientists and engineers]. Izdat. “Nauka”, Moscow, USSR, second revised edition, 1972. 400 pp. Translated from the English by V. L. Arlazarov, G. S. Razina and A. V. Uskov, Edited by R. S. Guter.

Khemming:1983:CTK

[Khe83] R. V. Khemming. Teoriya kodirovaniya i teoriya informatsii.

(Russian) [Coding theory and information theory]. “Radio i Svyaz′”, Moscow, USSR, 1983. 174 pp. Translated from the English by S. I. Gel′fand, Translation edited and with a preface by B. S. Tsybakov.

Kolpakov:2001:FAR

[KK01] Roman Kolpakov and Gregory Kucherov. Finding approximate repetitions under Hamming distance. Lecture Notes in Computer Science, 2161:170–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/2161/21610170.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/2161/
21610170.pdf.

Kolpakov:2003:FAR

[KK03] Roman Kolpakov and Gregory Kucherov. Finding approximate repetitions under Hamming distance. Theoretical Computer Science, 303(1):135–156, June 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:2015:FEE

[KKI^{+}15] Jaeyoung Kim, Myeongsu Kang, Md Shohidul Islam, Cheol-Hong Kim, and Jong-Myon Kim. A fast and energy-efficient Hamming decoder for software-defined radio using graphics processing units. The Journal of Supercomputing, 71(7): 2454–2472, July 2015. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1007/s11227-015-1396-x.

Klein:2013:CBH

[Kle13] Shmuel T. Klein. On the connection between Hamming codes, Heapsort and other methods. Information Processing Letters, 113(17): 617–620, August 30, 2013. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.com/
                                        science/article/pii/S0020019013001555.

Knoblauch:1988:BRW

[Kno88] M. Knoblauch. Buchbesprechung: R. W. Hamming: Information und Codierung, VCH Verlagsgesellschaft, Weinheim 1987. 270 Seiten, Preis: DM 74,-. Berichte der Bunsen-Gesellschaft fuer Physikalische Chemie, 1963– 1990, 92(8):943, August 1988. CO-DEN BBPCAX. ISSN 0005-9021.

Klein:2017:HPE

[KS17] Shmuel Klein and Dana Shapira. Hierarchical parallel evaluation of a Hamming code. Algorithms, 10(2): 50–??, April 2017.

Kim:2012:SSS

[KSSY12] Cheonshik Kim, Dongkyoo Shin, Dongil Shin, and Ching-Nung Yang. A (2, 2) secret sharing scheme based on Hamming code and AMBTC. Lecture Notes in Computer Science, 7197:129–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.com/chapter/
                                        10.1007/978-3-642-28490-8_14/.

Karadeniz:2015:NEB

[KY15] Suat Karadeniz and Bahattin Yildiz. New extremal binary self-dual codes of length 64 from R_{3}-lifts of the extended binary Hamming code. Designs, Codes, and Cryptography, 74(3):673–680, March 2015. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://link.

springer.com/article/10.1007/
s10623-013-9884-6.

Lange:2001:UEP

[LA01] Christoph Lange and Andreas Ahrens. On the undetected error probability for shortened Hamming codes on channels with memory. Lecture Notes in Computer Science, 2260:9–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/2260/22600009.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/2260/
22600009.pdf.

Liu:2008:RGH

[LCL08] Zihui Liu, Wende Chen, and Yuan Luo. The relative generalized Hamming weight of linear q-ary codes and their subcodes. Designs, Codes, and Cryptography, 48(2):111–123, August 2008. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://www.

springerlink.com/openurl.asp?
genre=article&issn=0925-1022&
volume=48&issue=2&spage=111.

Liu:2012:WIM

[LCWY12] Longcheng Liu, Yong Chen, Biao Wu, and Enyu Yao. Weighted inverse minimum cut problem under the sum-type Hamming distance. Lecture Notes in Computer Science, 7285:26–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.com/chapter/
                                        10.1007/978-3-642-29700-7_3/.

Lee:1998:RWH

[Lee98] Jan Lee. Richard Wesley Hamming: 1915–1998. IEEE Annals of the History of Computing, 20(2): 60–62, April/June 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:/

                                        /ieeexplore.ieee.org/iel4/85/
                                        14609/00667309.pdf.

Levenshtein:2004:SPC

[Lev04] V. I. Levenshtein. A system of polynomials for the complex Hamming spaces. In International Symposium on Information Theory, 2004. ISIT 2004. Proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004.

Lewis:2021:ICF

[Lew21] Harry R. Lewis. Ideas That Created the Future: Classic Papers of Computer Science. MIT Press, Cambridge, MA, USA, 2021. ISBN 0-262-04530-3. xxii + 495 pp. LCCN Q124.6-127.2.

Liu:2009:TAH

[Liu09] Andy Liu. Two applications of a Hamming code. College Mathematics Journal, 40(1):2–5, January 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

URL http://www.tandfonline.
com/doi/abs/10.1080/07468342.
2009.11922327.

Langtangen:2009:HE

[LL09] Hans Petter Langtangen and Olav Lysne. The Hamming experience. In Simula Research Laboratory, pages 33–35. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., October 2009.

Lin:2021:FCM

[LL22] Shu Lin and Juane Li. Fundamentals of Classical and Modern Errorcorrecting Codes. Cambridge University Press, Cambridge, UK, 2022. ISBN 1-009-06792-3 (e-pub), 1-316-51262-2 (hardcover). xxix + 812 pp. LCCN TK5102.96 .L53 2021.

Legg:2017:GVF

[LMWC17] Philip A. Legg, Eamonn Maguire, Simon Walton, and Min Chen. Glyph visualization: A fail-safe design scheme based on quasi-Hamming distances. IEEE Computer Graphics and Applications, 37(2):31–41, March/April 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

URL https://www.computer.org/
csdl/mags/cg/2017/02/mcg2017020031-
abs.html.

Loomis:2002:RWH

[LP02] Herschel H. Loomis and David S.

Potter. Richard W. Hamming (1915–1998). In Anonymous, editor, Memorial Tributes, volume 10, pages 121–124. National Academy Press, Washington, DC, USA, 2002. ISBN 0-309-08457-1, 0-585-14960-7 (e-book), 0-309-05575-X. LCCN TA139 .N34 2002. URL http://

                                        nap.edu/10403/chapter/23.

Liu:2014:LHZ

[LPH14] Xing Liu, Daiyuan Peng, and Hongyu Han. Low-hit-zone frequency hopping sequence sets with optimal partial Hamming correlation properties. Designs, Codes, and Cryptography, 73(1):167–176, October 2014. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://link.

                                        springer.com/article/10.1007/
                                        s10623-013-9817-4.

Long:2014:WHH

[LS14] Philip M. Long and Rocco A. Servedio. On the weight of halfspaces over Hamming balls. SIAM Journal on Discrete Mathematics, 28(3): 1035–1061, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liu:2010:RGH

[LWW10] Zihui Liu, Jie Wang, and Xin-Wen Wu. On the relative generalized Hamming weights of linear codes and their subcodes. SIAM Journal on Discrete Mathematics, 24(4): 1234–1241, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liu:2008:IMM

[LY08] Longcheng Liu and Enyu Yao. Inverse min–max spanning tree problem under the Weighted sum-type Hamming distance. Theoretical Computer Science, 396(1–3):28–34, May 10, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manz:2016:HGC

[Man16] Olaf Manz. Hamming- und Golay-Codes. In Fehlerkorrigierende Codes, pages 51–70. Springer Fachmedien Wiesbaden, December 2016.

Mandelberg:2018:RWH

[Man18] Martin Mandelberg. Richard Wesley Hamming: Man, Mathematician, Mentor. ????, ????,

2018.  ????  pp.  URL

https://richardwesleyhamming.
com/foreword-of-the-book. In

preparation, Fall 2018; see [Kel18b].

Mazurkiewicz:2003:SSH

[Maz03] Jacek Mazurkiewicz. Systolic simulation of Hamming neural network. In Neural Networks and Soft Computing, pages 867–872. Physica-Verlag HD, 2003.

Ma:2010:HCM

[MC10] Rick Ma and Samuel Cheng. Hamming coding for multiple sources. In 2010 IEEE International Symposium on Information Theory. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 2010.

McCall:1959:NMR

[MH59] D. W. McCall and R. W. Hamming.

Nuclear magnetic resonance in crystals. Acta Crystallographica, 12(2): 81–86, February 1959. CODEN AC-CRA9. ISSN 0365-110X (print), 1600-8642 (electronic). See [MH63].

McCall:1963:ENM

[MH63] D. W. McCall and R. W. Hamming. [Errata:] nuclear magnetic resonance in crystals. Acta Crystallographica, 16(10):1071, October 1963. CODEN ACCRA9. ISSN 0365-110X (print), 1600-8642 (electronic). See [MH59].

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Mickens:1985:MAP

[Mic85] Ronald E. Mickens, editor. Mathematical analysis of physical systems. Van Nostrand Reinhold Co., New York, NY, USA, 1985. ISBN 0-442-26077-6. x + 357 pp. LCCN QC20 .M36 1985.

Mitchell:2009:IHR

[Mit09] Gregory Mitchell. Investigation of Hamming, Reed–Solomon, and Turbo forward error correcting codes. Technical report, Defense Technical Information Center, ????, 2009.

Marinescu:2012:CQI

[MM12] Dan C. Marinescu and Gabriela M. Marinescu, editors. Classical and Quantum Information. Elsevier, Amsterdam, The Netherlands, 2012. ISBN 0-12-383874-6 (hardcover). xviii + 725 pp. LCCN QA76.889 .M363 2012.

Morgan:1998:RWH

[Mor98] Samuel P. Morgan. Richard Wesley Hamming (1915–1998). Notices of the American Mathematical Society, 45(8):972–977, September 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http:

//www.ams.org/notices/199808/
mem-morgan.pdf.

Morgan:2002:HRW

[Mor02] Samuel P. Morgan. Hamming, Richard Wesley (1915–1998), mathematician and computer scientist. In Mark C. (Mark Christopher) Carnes, editor, American national biography online: the life of a nation is told by the lives of its people, volume 9, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, July 2002. ISBN 0-19-520635-5 (24-volume set). LCCN CT213 .A68 1999.

Mittelmann:2010:EBQ

[MP10] Hans Mittelmann and Jiming Peng.

Estimating bounds for quadratic assignment problems associated with Hamming and Manhattan distance matrices based on semidefinite programming. SIAM Journal on Optimization, 20(6):3408–3426, ???? 2010. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (elec-

tronic). URL http://epubs.siam.
org/siopt/resource/1/sjope8/
v20/i6/p3408_s1.

Manthey:2005:ICH

[MR05] Bodo Manthey and Rüdiger Reischuk. The intractability of computing the Hamming distance. Theoretical Computer Science, 337(1–3): 331–346, June 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin:2002:DCP

[MS02] José Luis Fernández-Villacañas Martín and Mónica Sierra Sánchez. Does crossover probability depend on fitness and Hamming differences in genetic algorithms? Lecture Notes in Computer Science, 2415:389–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://link.springer-ny.
                                        com/link/service/series/0558/
                                        bibs/2415/24150389.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/2415/
                                        24150389.pdf.

Muniz:2006:IEH

[Mun06] Marcelo Muniz. Isometric embeddings of Z_{p}^{k} in the Hamming space F_{p}^{N} and Z_{p}^{k}-linear codes. Designs, Codes, and Cryptography, 41(2):147–152, November 2006. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?
                                        genre=article&issn=0925-1022&
                                        volume=41&issue=2&spage=147.

Morelos-Zaragoza:2002:AEC

[MZ02a] Robert H. Morelos-Zaragoza. The Art of Error Correcting Coding. John Wiley, New York, NY, USA, 2002. ISBN 0-471-49581-6, 0-470-84782-4 (e-book), 0-470-85247-X (e-book), 0-470-03569-2 (ebook). xvi + 221 pp. LCCN QA268 .M67 2002. URL http://

onlinelibrary.wiley.com/book/
10.1002/0470847824.

Morelos-Zaragoza:2002:HGR

[MZ02b] Robert H. Morelos-Zaragoza. Hamming, Golay and Reed–Muller codes. In The Art of Error Correcting Coding [MZ02a], chapter 2, pages 23–32. ISBN 0-471-49581-6, 0-470-84782-4 (ebook), 0-470-85247-X (e-book), 0-470-03569-2 (e-book). LCCN QA268 .M67 2002. URL http://

onlinelibrary.wiley.com/book/
10.1002/0470847824.

Olaya-Leon:2015:SGH

[OLGP15] Wilson Olaya-León and Claudia Granados-Pinzón. The second generalized Hamming weight of certain Castle codes. Designs, Codes, and Cryptography, 76(1):81– 87, July 2015. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://link.

springer.com/article/10.1007/
s10623-014-9981-1.

OConnor:2012:RWH

[OR12] J. J. O’Connor and E. F. Robertson. Richard Wesley Hamming. MacTutor History of Mathematics archive document., January 2012.

URL http://www-history.mcs.
st-andrews.ac.uk/Biographies/
Hamming.html.

Oren:2012:ASC

[ORSW12] Yossef Oren, Mathieu Renauld, François-Xavier Standaert, and Avishai Wool. Algebraic sidechannel attacks beyond the Hamming weight leakage model. Lecture Notes in Computer Science, 7428: 140–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/chapter/10.1007/
                                        978-3-642-33027-8_9/.

Ostergaard:2002:CSH

[Öst02] Patric R. J. Östergård. Classifying subspaces of Hamming spaces. Designs, Codes, and Cryptography, 27(3):297–305, December 2002. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (elec-

                                        tronic). URL http://ipsapp008.
                                        lwwonline.com/content/getfile/
                                        4630/44/9/abstract.htm;   http:
                                        //ipsapp008.lwwonline.com/content/
                                        getfile/4630/44/9/fulltext.pdf.

Ostergaard:2004:REF

[ÖV04] Patric R. J. Östergård and Alexander Vardy. Resolving the existence of full-rank tilings of binary Hamming spaces. SIAM Journal on Discrete Mathematics, 18(2):382–387, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.siam.
                                        org/sam-bin/dbq/article/43580.

Pang:1986:CCC

[PE86] King F. Pang and Abbas El Gamal. Communication complexity of computing the Hamming distance. SIAM Journal on Computing, 15 (4):932–947, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Perry:1993:RWH

[Per93] Tekla S. Perry. Richard W.

Hamming. IEEE Spectrum, 30 (5):80–82, May 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic). URL

http://www.readabstracts.com/
Engineering-and-manufacturing-
industries/Susan-Hackwood-Richard-
W-Hamming.html.

Phelps:1995:KNH

[PL95] Kevin T. Phelps and Mike Levan.

Kernels of nonlinear Hamming codes. Designs, Codes, and Cryptography, 6(3):247–257, November 1995. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (elec-

tronic). URL http://www.wkap.
nl/oasis.htm/93425.

Panek:2018:CHE

[PP18] Luciano Panek and Nayene Michele Paião Panek. Códigos de Hamming estendidos como códigos perfeitos. In ????, Proceeding Series of the Brazilian Society of Computational and Applied Mathematics. SBMAC: Sociedade de Matemática Aplicada e Computacional, Edifício Medical Center — Rua Maestro João Seppe, n. 900, 16. andar — Sala 163 — São Carlos/SP - CEP: 13561-120, Brazil, February 2018.

Parnas:1989:CRD

[PSvE^{+}89] David L. Parnas, W. L. Scherlis, M. H. van Emden, Jacques Cohen, R. W. Hamming, Richard M.

Karp, and Terry Winograd. Colleagues respond to Dijkstra’s comments. Communications of the ACM, 32(12):1405–1413, December 1989. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pirsic:2014:DFT

[PW14] Gottlieb Isabel Pirsic and Arne Winterhof. On discrete Fourier transform, ambiguity, and Hammingautocorrelation of pseudorandom sequences. Designs, Codes, and Cryptography, 73(2):319–328, November 2014. CODEN DC-CREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1007/s10623-013-9916-2.

Parisi:2006:HDB

[PZ06] Giorgio Parisi and Francesco Zamponi. On the high density behavior of Hamming codes with fixed minimum distance. Journal of Statistical Physics, 123(6):1145–1167, June 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

                                        springer.com/article/10.1007/
                                        s10955-006-9142-7.

Rahim:2017:BED

[Rah17] Robbi Rahim. Bit error detection and correction with Hamming code algorithm. ????, September 2017.

Rigas:2015:EMV

[REF15] Ioannis Rigas, George Economou, and Spiros Fotopoulos. Efficient modeling of visual saliency based on local sparse representation and the use of Hamming distance. Computer Vision and Image Understanding: CVIU, 134 (??):33–45, May 2015. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.sciencedirect.com/
science/article/pii/S1077314215000223.

Ren:2014:NSF

[RFZ14] Wenli Ren, Fang-Wei Fu, and Zhengchun Zhou. New sets of frequency-hopping sequences with optimal Hamming correlation. Designs, Codes, and Cryptography, 72 (2):423–434, August 2014. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http:

//link.springer.com/article/
10.1007/s10623-012-9774-3.

Rurik:2016:HCE

[RM16] William Rurik and Arya Mazumdar. Hamming codes as errorreducing codes. In 2016 IEEE Information Theory Workshop (ITW). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 2016.

Robertson:1987:EDC

[Rob87] James E. Robertson. Errordetection and correction for addition and subtraction through use of higher radix extensions of Hamming codes. In Irwin and Stefanelli [IS87], pages 226–229. ISBN 0-8186-0774-2 (paperback), 0-8186-4774-4 (microfiche), 0-8186-8774-6 (case). LCCN QA 76.9 C62 S95 1987. URL http://

www.acsel-lab.com/arithmetic/

                                        arith8/papers/ARITH8_Robertson.
                                        pdf.

Robinson:2001:WMN

[Rob01] Sara Robinson. Why mathematicians now care about their hat color. New York Times, ??(??):F5, April 10, 2001. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X,

                                        1553-8095. URL https://search.
                                        proquest.com/hnpnewyorktimes/
                                        docview/92030606/.

Rifa:2012:IHC

[RSV12] J. Rifà, F. I. Solov’eva, and M. Villanueva. Intersection of Hamming codes avoiding Hamming subcodes. Designs, Codes, and Cryptography, 62(2):209–223, February 2012. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?
                                        genre=article&issn=0925-1022&
                                        volume=62&issue=2&spage=209.

Rifa:2015:EIH

[RSV15a] J. Rifà, F. I. Solov’eva, and M. Villanueva. Erratum to: Intersection of Hamming codes avoiding Hamming subcodes. Designs, Codes, and Cryptography, 74(1):283, January 2015. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://link.

                                        springer.com/article/10.1007/
                                        s10623-014-0011-0;   http://
                                        link.springer.com/content/pdf/
                                        10.1007/s10623-014-0011-0.pdf.

Rifa:2015:SEH

[RSV15b] Josep Rifà, Faina I. Solov’eva, and Mercè Villanueva. Self-embeddings of Hamming Steiner triple systems of small order and APN permutations. Designs, Codes, and Cryptography, 75(3):405–427, June 2015. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://link.

springer.com/article/10.1007/
s10623-013-9909-1.

Ramezanpour:2012:CAS

[RZ12] A. Ramezanpour and R. Zecchina.

Cavity approach to sphere packing in Hamming space. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 85(2):021106, February 6, 2012. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (elec-

tronic).  URL https://link.
aps.org/doi/10.1103/PhysRevE.
85.021106.

S:1975:BRR

[S.75] J. S. Book review: R. W. Hamming, Numerical Methods for Scientists and Engineers, 2nd ed., McGraw-Hill Book Co., New York, 1973, ix + 721 pp., 24 cm. Mathematics of Computation, 29(130):648– 649, May 1975. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/
mcom/1975-29-130/S0025-5718-
75-99678-7/S0025-5718-75-99678-
7.pdf.

Sack:2016:RHH

[Sac16] Harald Sack. Richard Hamming and the Hamming code. Web blog., February 11, 2016. URL http:

//scihi.org/richard-hamming-
hamming-code/.

Salman:2013:HDD

[Sal13] Raied Salman. Hamming distance and data compression of 1-D CA. In Computer Science & Information Technology (CS & IT). Academy & Industry Research Collaboration Center (AIRCC), ????, May 2013.

Samorodnitsky:2017:IFH

[Sam17] Alex Samorodnitsky. An inequality for functions on the Hamming cube. Combinatorics, Probability and Computing, 26(3):468–480, May 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL /core/product/

                                        3770E7921E125EF4A73A413E396348CB.

Scaife:1974:SNA

[Sca74] B. K. P. (Brendan Kevin Patrick) Scaife, editor. Studies in numerical analysis: papers in honour of Cornelius Lanczos. Academic Press, New York, USA, 1974. ISBN 0-12-621150-7. xxii + 333 pp. LCCN QA297

                                        .S86.  URL http://catalog.
                                        hathitrust.org/Record/000575557.

Published for the Royal Irish Academy.

Schulz:1991:HCE

[Sch91] Ralph-Hardo Schulz. Hamming-Codes und erweiterte Hamming-Codes. (German) [Hamming codes and extended Hamming codes]. In Codierungstheorie. (German) [Coding theory], pages 111–115. Vieweg + Teubner Verlag, Braunschweig, Germany, 1991.

Schulz:2003:HCE

[Sch03] Ralph-Hardo Schulz. Hamming-Codes und erweiterte Hamming-Codes. (German) [Hamming codes and extended Hamming codes]. In Codierungstheorie. (German) [Coding theory], pages 112–116. Vieweg + Teubner Verlag, Braunschweig, Germany, 2003.

Serov:2011:LDH

[Ser11] A. A. Serov. Limit distribution of the Hamming distance from the random Boolean function to the set of affine functions. Theory of Probability and its Applications, 55 (4):717–722, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:/

/epubs.siam.org/tvp/resource/
1/tprbau/v55/i4/p717_s1.

Stearns:1990:DSA

[SH90] Samuel D. Stearns and Don R.

Hush. Digital Signal Analysis. Prentice Hall signal processing series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1990. ISBN 0-13-213117-X, 0-13-211772-X (paperback). xviii + 440 pp. LCCN TK5102.5 .S698 1990.

Stearns:1999:DVA

[SH99] Samuel D. Stearns and Don R.

Hush, editors. Digitale Verarbeitung analoger Signale. (German) [Digital Analysis of Analog Signal]. Walter de Gruyter, New York, NY, USA and Berlin, Germany, December 1999. 571 pp.

Shallit:2009:HDC

[Sha09] Jeffrey Shallit. Hamming distance for conjugates. Discrete Mathematics, 309(12):4197–4199, June 2009. CODEN DSMHA4. ISSN 0012-365X (print), 1872-681X (electronic).

Shepherd:1995:PPH

[She95] S. J. Shepherd. Primitive polynomials over GF(2) of Hamming weight 3 and 5 up to high order. Report 573, Electronic and Electrical Engineering Department, University of Bradford, Bradford, Yorkshire, UK, March 1995.

Sivakoff:2014:SPD

[Siv14] David Sivakoff. Site percolation on the d-dimensional Hamming torus. Combinatorics, Probability and Computing, 23(2):290– 315, March 2014. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Sarvepalli:2010:DQC

[SK10a] Pradeep Sarvepalli and Andreas Klappenecker. Degenerate quantum codes and the quantum Hamming bound. Physical Review A (Atomic, Molecular, and Optical Physics), 81(3):032318, March 19, 2010. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446,

                                        1538-4519. URL https://link.
                                        aps.org/doi/10.1103/PhysRevA.
                                        81.032318.

Savari:2010:WHM

[SK10b] S. A. Savari and J. Kliewer. When Huffman meets Hamming: a class of optimal variable-length error correcting codes. In Storer and Marcellin [SM10], pages 327–336. ISBN 0-7695-3994-

7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5453476.

Storer:2010:DPD

[SM10] James A. (James Andrew) Storer and Michael W. Marcellin, editors. DDC 2010: proceedings: Data Compression Conference: 24–26 March 2010, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5453521.

Solomon:2000:GED

[Sol00] G. Solomon. Golay encoding/ decoding via BCH–Hamming. Computers and Mathematics with Applications, 39(11):103–108, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0898122100001140.

Sakai:2002:AES

[SS02] Yasuyuki Sakai and Kouichi Sakurai. Algorithms for efficient simultaneous elliptic scalar multiplication with reduced joint Hamming weight representation of scalars. Lecture Notes in Computer Science, 2433:484–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/2433/24330484.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/2433/
24330484.pdf.

Stakhov:2018:MCS

[Sta18] Alexey Stakhov. Mission-critical systems, paradox of Hamming code, row hammer effect, ‘Trojan Horse’ of the binary system and numeral systems with irrational bases. The Computer Journal, 61(7):1038–1063, July 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://academic.
oup.com/comjnl/article/61/7/
1038/4430323.

Stinson:2002:SBS

[Sti02] D. R. Stinson. Some babystep giant-step algorithms for the low Hamming weight discrete logarithm problem. Mathematics of Computation, 71(237):379–391, January 2002. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:/

/www.ams.org/journal-getitem?
pii=S0025-5718-01-01310-2;
http://www.ams.org/mcom/2002-
71-237/S0025-5718-01-01310-2/
S0025-5718-01-01310-2.dvi;
http://www.ams.org/mcom/2002-
71-237/S0025-5718-01-01310-2/
S0025-5718-01-01310-2.pdf;
http://www.ams.org/mcom/2002-
71-237/S0025-5718-01-01310-2/
S0025-5718-01-01310-2.ps;   http:
//www.ams.org/mcom/2002-71-237/
S0025-5718-01-01310-2/S0025-
5718-01-01310-2.tex.

Swartzlander:1990:CAb

[Swa90] Earl E. Swartzlander, Jr. Computer Arithmetic, volume 2. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-8945-5. ix + 396 pp. LCCN QA76.9 .C62C66 1990. This is part of a two-volume collection of influential papers on the design of computer arithmetic. See also [?].

Swartzlander:2015:CAa

[Swa15a] Earl E. Swartzlander, Jr., editor.

Computer Arithmetic, volume 1. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2015. ISBN 981-4651-56-7 (vol. 1; hardcover), 981-4651-57-5, 981-4641-47-2 (e-book). ???? pp. LCCN QA76.6 .C633 2015 vol. 1.

Swartzlander:2015:CAb

[Swa15b] Earl E. Swartzlander, Jr., editor.

Computer Arithmetic, volume 2. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2015. ISBN 981-4641-46-4 (hardcover), 981-4641-47-2 (e-book). xxxviii + 446 pp. LCCN QA76.6 .C633 2015 vol. 2.

Swartzlander:2015:CAc

[Swa15c] Earl E. Swartzlander, Jr., editor.

Computer Arithmetic, volume 3. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2015. ISBN 981-4651-13-3 (hardcover), 981-4641-47-2 (e-book). xvii + 451 pp. LCCN QA76.6 .C633 2015 vol. 3.

Swetina:1989:FSM

[Swe89] J. Swetina. First and second moments and the mean Hamming distance in a stochastic replicationmutation model for biological macromolecules. Journal of Mathematical Biology, 27(4):463–483, August 1989. CODEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic). URL http://link.

springer.com/article/10.1007/
BF00290640.

Sutton:2012:CMB

[SWH12] Andrew M. Sutton, L. Darrell Whitley, and Adele E. Howe. Computing the moments of kbounded pseudo-Boolean functions over Hamming spheres of arbitrary radius in polynomial time. Theoretical Computer Science, 425(1): 58–74, March 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/S0304397511001101.

Takashima:1996:HWT

[Tak96] Keizo Takashima. On Hamming weight test and sojourn time test of m-sequences. Monte Carlo Methods and Applications, 2(4): 331–340, ???? 1996. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/
view/j/mcma.1996.2.issue-4/mcma.
1996.2.4.331/mcma.1996.2.4.331.
xml.

Terrell:1978:BRD

[Ter78] T. J. Terrell. Book review: Digital Filters Hamming, Richard W. (Prentice-Hall Inc., 1977, 226 pp., £14.35). International Journal of Electrical Engineering Education, 15(3):288, July 1978. CODEN IJEEAF. ISSN 0020-

7209. URL http://journals.
sagepub.com/doi/pdf/10.1177/
002072097801500327.

Terrell:1984:BRD

[Ter84] T. J. Terrell. Book review: Digital Filters, (Second Edition): Hamming, R. W. (Prentice-Hall, 1983, 257 pp., £26.05). International Journal of Electrical Engineering Education, 21(1):94–95, January 1984. CODEN IJEEAF. ISSN 0020-7209.

Tukey:1949:MNC

[TH49] John W. Tukey and R. W. Hamming. Measuring noise color. Technical memorandum, Bell Laboratories, Murray Hill, NJ, USA, 1949. 1–127 pp.

Thompson:1973:BRC

[Tho73] Michael Thompson. Book review:

Computers and Society, by Richard W. Hamming. Leonardo (Oxford, England), 6(4):368, Autumn 1973. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282 (elec-

tronic). URL https://www.jstor.
org/stable/1572868.

Trevisan:1997:WHM

[Tre97] Luca Trevisan. When Hamming meets Euclid: the approximability of geometric TSP and MST (extended abstract). In ACM [ACM97], pages 21– 29. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997. URL http:

//www.acm.org/pubs/articles/
proceedings/stoc/258533/p21-
trevisan/p21-trevisan.pdf;
http://www.acm.org/pubs/citations/
proceedings/stoc/258533/p21-

trevisan/. ACM order no. 508970.

Trevisan:2000:WHM

[Tre00] Luca Trevisan. When Hamming meets Euclid: The approximability of geometric TSP and Steiner tree. SIAM Journal on Computing, 30(2):475–485, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/dbq/
article/35273.

Tulpan:2014:TPP

[TSM14] Dan Tulpan, Derek H. Smith, and Roberto Montemanni. Thermodynamic post-processing versus GCcontent pre-processing for DNA codes satisfying the Hamming distance and reverse-complement constraints. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 11(2):441–452, March 2014. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Tannenbaum:2004:IDL

[TSS04] Emmanuel Tannenbaum, James L. Sherley, and Eugene I. Shakhnovich. Imperfect DNA lesion repair in the semiconservative quasispecies model: Derivation of the Hamming class equations and solution of the single-fitness peak landscape. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 70 (6):061915, December 30, 2004. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic).

URL https://link.aps.org/doi/
10.1103/PhysRevE.70.061915.

Tromp:1995:SWB

[TZZ95] J. Tromp, L. Zhang, and Y. Zhao.

Small weight bases for Hamming codes. Lecture Notes in Computer Science, 959:235–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tromp:1997:SWB

[TZZ97] John Tromp, Louxin Zhang, and Ying Zhao. Small weight bases for Hamming codes. Theoretical Computer Science, 181(2):337–345, July 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/
browse.cgi?year=1997&volume=
181&issue=2&aid=2480.

vanderHofstad:2010:RSH

[vdHL10] Remco van der Hofstad and Malwina J. Luczak. Random subgraphs of the 2D Hamming graph: the supercritical phase. Probability Theory and Related Fields, 147(1–2):1– 41, May 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/
s00440-009-0200-3.

Viswanath:2006:MCL

[VR06] G. Viswanath and B. Sundar Rajan. Matrix characterization of linear codes with arbitrary Hamming weight hierarchy. Linear Algebra and Its Applications, 412(2–3):396– 407, January 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanderGeer:1994:GHW

[VV94] G. Van der Geer and M. Van der Vlugt. Generalized Hamming weights of Melas codes and dual Melas codes. SIAM Journal on Discrete Mathematics, 7(4):554–559, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Warren:2013:HD

[War13] Henry S. Warren. Hacker’s Delight. Addison-Wesley, Reading, MA, USA, second edition, 2013. ISBN 0-321-84268-5 (hardcover). xvi + 494 pp. LCCN QA76.6 .W375 2013. URL http://www.

                                        pearsonhighered.com/educator/
                                        product/Hackers-Delight/9780321842688.
                                        page.

Wong:2014:PHD

[WK14] Kok-Seng Wong and Myung Ho Kim. On private Hamming distance computation. The Journal of Supercomputing, 69(3):1123– 1138, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1007/s11227-013-1063-z.

Wang:2014:HEK

[WZNM14] Feng Wang, Wan-Lei Zhao, Chong-Wah Ngo, and Bernard Merialdo. A Hamming embedding kernel with informative bagof-visual words for video semantic indexing. ACM Transactions on Multimedia Computing, Communications, and Applications, 10 (3):26:1–26:??, April 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2017:NBP

[WZYG17] Xin Wang, Yiwei Zhang, Yiting Yang, and Gennian Ge. New bounds of permutation codes under Hamming metric and Kendall’s τ-metric. Designs, Codes, and Cryptography, 85(3):533–545, December 2017. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL https://link.

springer.com/article/10.1007/
s10623-016-0321-5.

Zhao:2013:EHW

[ZGZ^{+}13] Xinjie Zhao, Shize Guo, Fan Zhang, Tao Wang, Zhijie Shi, Huiying Liu, Keke Ji, and Jing Huang. Efficient Hamming weightbased side-channel cube attacks on PRESENT. The Journal of Systems and Software, 86(3):728– 743, March 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/S0164121212003081.

Ziegler:2001:CHG

[Zie01] Günter M. Ziegler. Coloring Hamming graphs, optimal binary codes, and the 0/1-Borsuk problem in low dimensions. Lecture Notes in Computer Science, 2122:159–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://link.springer-ny.
                                        com/link/service/series/0558/
                                        bibs/2122/21220159.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/2122/
                                        21220159.pdf.

Zhuang:2013:CCE

[ZLD13] Zhuojun Zhuang, Yuan Luo, and Bin Dai. Code constructions and existence bounds for relative generalized Hamming weight. Designs, Codes, and Cryptography, 69(3): 275–297, December 2013. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1007/s10623-012-9657-7.

Zhou:2012:AHC

[ZTYP12] Zhengchun Zhou, Xiaohu Tang, Yang Yang, and Udaya Parampalli. On the aperiodic Hamming correlation of frequency-hopping sequences from norm functions. Lecture Notes in Computer Science, 7280:148– 158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/chapter/10.1007/
                                        978-3-642-30615-0_14/.